### v1

# Low-Luminosity Type IIP Supernovae from the Zwicky Transient Facility Census of the Local Universe. II: Lightcurve Analysis

Abstract
The Zwicky Transient Facility Census of the Local Universe survey yielded a sample of 330 Type IIP supernovae (SNe) with well-constrained peak luminosities. In paper I (Das et al., 2025), we measured their luminosity function and volumetric rate. Here (paper II), we present the largest systematic study of lightcurve properties for Type IIP SNe from a volume-limited survey, analyzing a selected subset of 129 events, including 16 low-luminosity Type IIP (LLIIP) SNe with M mag. We find that plateau slope correlates with peak brightness, with many LLIIP SNe showing positive slopes—suggesting smaller progenitor radii and distinct density profiles compared to brighter Type IIP SNe. The plateau duration shows only a weak dependence on peak brightness, likely suggesting binary interaction. One SN exhibits a plateau-to-tail drop of mag, consistent with an electron-capture or failed SN with very low or zero nickel mass. We derive explosion and progenitor parameters of the entire Type IIP SN sample using semi-analytical and radiation-hydrodynamical models. Based on radiation-hydrodynamical model fitting, LLIIP SNe are characterized by low nickel masses ( ), low explosion energies ( erg), low ejecta masses ( ), and ZAMS masses below . In comparison, the full Type IIP SN sample spans a wider range with nickel masses (– ), explosion energies (–erg), ejecta masses ( ), and ZAMS masses (– ). We find strong correlations between peak brightness, explosion energy, and nickel mass that extend to the low-luminosity end. We conclude that LLIIP SNe represent the faint, low-energy end of the Type IIP population and originate from the lowest-mass core-collapse progenitors.
1 Introduction
Core-collapse supernovae (CCSNe) are the explosive deaths of massive stars that play crucial roles in galactic chemical evolution, the formation of new stars, and the creation of compact neutron stars and black holes. Moreover, their luminosities, rates, and association with massive star formation make them essential tools for probing key astrophysical parameters across vast cosmic volumes. The most common type of core-collapse SNe are Type II SNe (e.g., Li et al., 2011; Shivvers et al., 2017). They arise from the explosions of stars that retain significant portions of their hydrogen envelopes.
Low-luminosity Type IIP supernovae (LLIIP SNe), defined here as those with mag, as motivated by prior literature (e.g., Spiro et al., 2014), are particularly intriguing because they are thought to originate from progenitor stars with just enough mass to undergo core collapse. These events likely arise from 8–12 progenitors, as supported by pre-SN progenitor imaging (Maund et al., 2005; Li et al., 2006; Mattila et al., 2008; O’Neill et al., 2019; Van Dyk et al., 2023), light curve simulations (Pumo et al., 2017; Fraser et al., 2011), evolutionary models of low-mass red supergiants (RSGs) (Lisakov et al., 2017), and nebular spectroscopy (Jerkstrand et al., 2018). They provide valuable insight into the lower mass threshold for core collapse, which defines the boundary between stars that form white dwarfs and those that form neutron stars or black holes. LLIIP SNe also serve as key candidates for probing the upper mass limit of super-asymptotic giant branch (sAGB) star formation, exploring whether they can explode as electron-capture supernovae (ECSN) (Nomoto, 1984), and understanding their role in shaping elemental abundances in subsequent generations of stars (Tsujimoto & Shigeyama, 1999; Siess & Pumo, 2006). Furthermore, a typical initial mass function (IMF) predicts that more than 40% of CCSN progenitors fall within the 8–12 range, with 25% within the 8–10 subset (Sukhbold et al., 2016). It has also been suggested that LLIIP SNe may instead originate from more massive RSGs with significant fallback (e.g., Zampieri et al., 2003). Thus, systematic studies of LLIIP SNe are essential for improving our understanding of the progenitors, explosion properties, and the dust and chemical enrichment contributed by nearly half of all massive stars that undergo core-collapse.
However, to date, no study of the lightcurve properties of LLIIP SNe has been conducted using data from a systematic SN survey. Due to their low luminosity, they are relatively difficult to detect and classify. There are about a dozen such objects presented in the literature (Turatto et al., 1998; Spiro et al., 2014; Jäger et al., 2020; Müller-Bravo et al., 2020; Reguitti et al., 2021; Yang et al., 2021; Valerin et al., 2022; Kozyreva et al., 2022; Bostroem et al., 2023; Teja et al., 2024; Dastidar et al., 2025). In Part I of the series of three papers on LLIIP SNe (Das et al., 2025), we measured the luminosity function and volumetric rate of 36 LLIIP SNe (and 320 Type IIP SNe) from the Zwicky Transient Facility Census of the Local Universe (ZTF CLU; De et al., 2020) survey. This paper delves into their lightcurve properties.
One of the most used methods to analyze the progenitor and explosion properties is modeling of SN lightcurves. It is well known that lightcurves are sensitive to the physical properties of their progenitors (final masses and radii), as well as the properties of the explosion itself (released energy, amount of synthesized radioactive nickel and its distribution). In this paper, we analyze the observed properties of the multi-band lightcurves of 16 LLIIP SNe. We also infer physical explosion parameters using semi-analytical and radiation-hydrodynamical models that link SN observables to progenitor properties (e.g., Litvinova & Nadezhin, 1983; Hamuy, 2003; Kasen & Woosley, 2009; Goldberg et al., 2019; Morozova et al., 2018; Moriya et al., 2023). Such models have been previously applied to both individual events and larger samples of Type II SNe (Hamuy, 2003; Arcavi et al., 2011; Anderson et al., 2014; Spiro et al., 2014; Sanders et al., 2015; Valenti et al., 2016; Gutiérrez et al., 2017; Morozova et al., 2018; Martinez et al., 2022a, b; Subrayan et al., 2023; Silva-Farfán et al., 2024; Fang et al., 2025). To place these results in a broader context, we also present the lightcurve analysis of the entire sample of 129 Type IIP SNe, obtained from the largest systematic volume-limited SN survey to date.
This paper is structured as follows: in Section 2, we define the sample selection criteria. The data and extinction correction methods are described in Sections 3 and 4, respectively. Sections 5 and 6 detail the fitting methods and results. We discuss the implications of the derived physical properties in Section 7 and conclude in Section 8.
2 Sample Selection
We refer to Paper I (Das et al., 2025) for the complete sample selection criteria to select 330 Type IIP SNe from the ZTF (Bellm et al., 2019; Graham et al., 2019; Dekany et al., 2020; Masci et al., 2019) CLU experiment. To select the sample for lightcurve analysis, we require that the end of the plateau phase is observed in the -band lightcurve. This is to ensure we can extract reliable progenitor and explosion parameters from the lightcurves. The epoch of the end of the plateau () is determined as the epoch where the derivative of the slope of the -band lightcurve () reach a local minimum, and its value is mag day-2, indicating a sharp transition (see Section 5.1 and Figure 1 for details). This marks the end of the plateau phase in the lightcurve evolution. Using this criterion, we identified 129 candidates that fulfill this selection criterion. Out of these, 68 objects have -band data points in the radioactive tail phase, while 61 SNe do not have an observed radioactive tail.
3 Data
In this section, we describe the photometric and spectroscopic data used.
3.1 Optical photometry
We perform forced photometry on the ZTF difference images using the ZTF forced-photometry service developed by Masci et al. (2019) in , and bands. For this work, we consider anything less than a 3 detection an upper limit. In the tail phase, for a few SNe we also took photometry in the , , bands with the Optical Imager (IO:O) at the 2.0 m robotic Liverpool Telescope (Steele et al., 2004) and the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (Djupvik & Andersen, 2010). The P60 and LT data sets were processed with the FPipe image subtraction pipeline (Fremling et al., 2016), using Sloan Digital Sky Survey (SDSS; Ahn et al., 2012) and PanSTARRS (PS1; Chambers et al., 2016) reference frames. See Das et al. (2025) for details on photometric observations and reductions. Figure 1 shows a few representative lightcurves as examples. The photometry data and lightcurve plots for all the SNe in the sample will be made available on Zenodo.
3.2 Optical spectroscopy
For each transient, at least one spectrum is usually obtained close to the peak luminosity to establish an initial spectroscopic classification. We primarily utilize the Double Beam Spectrograph (DBSP; Oke & Gunn, 1982) on the Palomar 200-inch Hale telescope and the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al., 2018; Rigault et al., 2019; Kim et al., 2022) on the Palomar 60-inch telescope for this purpose. Other instruments used to obtain classification spectra include: the Alhambra Faint Object Spectrograph and Camera at the Nordic Optical Telescope (NOT; Djupvik & Andersen, 2010), the Spectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al., 2014) on the Liverpool Telescope, the Low-Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I telescope. See Das et al. (2025) for details on spectral observations and reductions.
We employ the SuperNova Identification (SNID; Blondin & Tonry, 2007) code for these classifications. For spectra with significant host galaxy contamination, we instead utilize superfit (Howell et al., 2005).
4 Extinction Correction
We corrected for Galactic extinction using the maps provided by Schlafly & Finkbeiner (2011). For the corrections, we use the extinction law described by Cardelli et al. (1989) with a value of .
To estimate the host-galaxy extinction, we use the average color of the sample and measure the host assuming that the SN color exceeding the 1 template is caused by host extinction. See Das et al. (2025) for the intrinsic color template and further details.
5 Analysis
5.1 Observed lightcurve properties
Parameter | Phase / Slope condition | condition |
Phase 60 days; earlier of slope 0.01 mag day-1 or max. | Max. | |
Phase days; slope mag day-1 | Min. mag day-2 | |
Phase and days; slope 0 | Max. mag day-2 | |
Time between half-max. brightness ( mag) and peak epoch | Not applicable |
We employed a Gaussian Process (GP) algorithm111https://george.readthedocs.io/ (Ambikasaran et al., 2015) to model the -band and -band lightcurves. The top panel of Figure 1 displays the interpolated -band lightcurve, the middle panel shows its first derivative with respect to time (i.e., the slope, in mag day-1), and the bottom panel shows the time derivative of the slope, denoted as (in mag day-2).
The -band lightcurve parameters are measured using the following criteria:
Plateau onset (): Defined as the earlier of (i) the epoch when the slope becomes shallower than 0.01 mag day-1, or (ii) the epoch when the change in slope with respect to time () reaches its maximum value, measured within 60 days after first detection.
Plateau end (): The epoch of plateau end () is identified after 40 days, when the lightcurve begins to steeply decline. This point is determined as the epoch where reaches a minimum below mag day-2, indicating a sharp transition.
Tail onset (): Defined as the epoch post-plateau when reaches a maximum exceeding mag day-2, provided this occurs within 50 days after the plateau ends.
Rise time (): Defined as the time interval between the epoch when the -band magnitude is 0.75 mag fainter than peak (half-maximum) and the epoch of maximum light.
To estimate uncertainties in these observables, we adopt a Monte Carlo approach. We simulate 50 realizations of each lightcurve based on Gaussian noise scaled to the flux errors. For each realization, we recompute the observables using the same criteria. The standard deviation of the resulting distribution of values is used as the uncertainty on the measurements of these observables. The plateau duration is measured as . The optically thick plateau duration (OPTd) is measured as the duration between the explosion epoch and the midpoint of the plateau end and tail onset epochs. To identify ECSN candidates and failed SNe with a sharp drop from the plateau and no nickel mass, we measure the drop in -band magnitude from the plateau end to the first epoch of tail onset when it exists. If a tail does not exist, this is marked as “Limit = yes” in Table A or indicated as a V-shaped triangle in Figure 3. The criteria are summarized in Table 1.
5.2 Measuring photospheric velocities
We derive expansion velocities of the H I line by measuring the minima of the absorption troughs in P-Cygni profiles. A third-order polynomial is fit to the region surrounding the absorption minimum to estimate the velocity. For spectra with low resolution or strong host galaxy contamination, we manually inspect and identify the minimum. Uncertainties are estimated via a Monte Carlo method. We first compute a noise spectrum by subtracting a heavily smoothed version of the original spectrum. The standard deviation of this residual defines the spectral noise. We then generate perturbed realizations by adding Gaussian noise of equal standard deviation to the smoothed spectrum and remeasure the velocity for each realization. The 1 spread in the resulting velocity distribution is adopted as the measurement uncertainty. Only measurements with a standard deviation less than 50% of their velocity are considered reliable and included. The measured velocities and their errors are documented in Table F.
We find that LLIIP SNe exhibit systematically lower H I velocities compared to their higher-luminosity counterparts at all phases. To quantify this trend, we perform a linear fit to the log of the velocity as a function of phase. The best-fit relations are shown in Figure 6. The best-fit for the LLIIP sample is: , and for the full sample: , where is the expansion velocity in km s-1 and is the phase in days since explosion. While both LLIIP SNe and brighter IIP SNe show comparable decline rates, LLIIP SNe exhibit lower velocities. At 50 days post-explosion, the average velocity inferred from these fits is approximately km s-1 for LLIIP SNe and km s-1 for SNe with mag.
5.3 Bolometric Luminosity
First, we measure the bolometric luminosity of the SNe in our sample. Since we only have the - and -band magnitudes available for most of the SNe in our sample, we estimate the bolometric luminosities using the bolometric correction (BC) approach from Lyman et al. (2014) to estimate the bolometric magnitudes:
, where is the bolometric magnitude, and is the absolute magnitude of the SN in -band filter. Jäger et al. (2020) and Yang et al. (2021) showed that the bolometric luminosity of LLIIP SNe measured by bolometric correction approach from Lyman et al. (2014) is similar to that obtained by fitting a blackbody function to multiband photometry, which supports the validity of the empirical correlation for the bolometric correction given by Lyman et al. (2014).
We bin the data into intervals of 2 days and use - and -band photometry to get the bolometric luminosity for that epoch. If only one band of photometry is available in a bin, then we use the color from the nearest epoch where photometry data for both bands is available within 10 days to calculate the bolometric luminosity.
The distribution of the plateau bolometric luminosity is shown in the left panel of Figure 4. The full sample has a median peak luminosity of . LLIIP SNe have lower peak luminosities with a median of . We find a strong correlation between the mid-plateau bolometric luminosity () and the nickel mass (Figure 4). The nickel mass is measured by fitting the radioactive tail (see Section 5.4.1). The best-fit relation is
with a Pearson correlation coefficient of and a -value .
5.4 Estimating physical and explosion parameters
In order to estimate progenitor and explosion parameters of the Type IIP sample, we use semi-analytic models Nagy & Vinkó (2016) and radiation hydrodynamical models from Moriya et al. (2023). We also use scaling relations from Goldberg et al. (2019) to constrain the explosion parameters of LLIIP SNe.
5.4.1 Lightcurve fitting to Semi-analytical models
The model in Nagy & Vinkó (2016) is based on a two-component configuration consisting of a uniform dense stellar core and an extended low-mass envelope where the density decreases as an exponential function. Nagy & Vinkó (2016); Jäger et al. (2020) claim that the results from the two-component semi-analytic LC model are consistent with current state-of-the-art calculations for Type II SNe. They are useful to derive estimates or constraints of basic parameters like the explosion energy, ejected mass, and initial radius of the progenitor as well as the amount of synthesized radioactive nickel. These models assume spherical symmetry and do not adequately model the initial transient behavior at early stages (t20 days). Thus, we only fit for lightcurve data 20 days after the explosion. We use the MCMC routine in Jäger et al. (2020) to fit these semi-analytical expressions. The Python code used for these fits will be made available on GitHub after publication. The best-fit lightcurves for the SNe with and without radioactive tails are shown in Figure 7. The nickel mass is measured from the lightcurve tail by fitting for both the nickel mass and the gamma-ray trapping efficiency parameter as free parameters. The priors used for the fits are shown in Table B. We fit the semi-analytic models to Type IIP SNe. After marginalization, we have estimates with confidence intervals (1) for each of these parameters. The distribution of the best-fit parameters and their correlations are shown in Figure 9 and Figure 13. The best-fit values for each SN are listed in Table D. The sample statistics are summarized in Table 3.
5.4.2 Lightcurve fitting to radiation-hydrodynamical Type II model grids
Moriya et al. (2023) presented a comprehensive model grid containing 228,016 synthetic Type II SN light curves, based on calculations using the radiation-hydrodynamical code STELLA. This set of model grids has previously been used in Subrayan et al. (2023) and Silva-Farfán et al. (2024) to infer explosion parameters for a sample of 45 and 186 Type II SNe, respectively. Here we also include previously unpublished models for lower-mass progenitors ( ) with low explosion energies (1.0, 2.0, 3.0, 4.0, ). The overall parameter space covered by the model grid is progenitor masses (9, 10, , , , and at the zero-age main sequence (ZAMS), solar metallicity), explosion energies (, , , , , , , , , and ), nickel masses (, , , , , , , , and ), mass-loss rates (, , , , , , , , and yr-1 with a wind velocity of 10 km s-1), circumstellar matter radii ( and ), and ten circumstellar structures ( and ), where is a wind structure parameter determined by the efficiency of wind acceleration (Moriya et al., 2023). The radioactive nickel is assumed to be uniformly mixed up to half the mass of the hydrogen-rich envelope.
We use this model grid to measure the explosion parameters of the CLU Type IIP SNe. Using Bayesian Inference methods requires the models to be finely sampled within the parameter space. However, because the model grid is discrete and non-uniform, we extrapolate the model parameter by doing a weighted average of the model magnitudes weighted by the separation of the vector space of the nearest-neighbour models from the model vector to be sampled. For a given parameter vector , the method finds the closest models and weighs them appropriately using
where is the magnitude for a given at time , where
5.4.3 Scaling Relations
Scaling relations have been commonly used to infer explosion parameters from lightcurve observables (Popov, 1993; Kasen & Woosley, 2009). Ejecta mass (), progenitor radius (), and explosion energy () are commonly used as the primary independent variables in scaling relations. In this paper, we use scaling relations derived by Goldberg et al. (2019). We use the following relations to constrain the ejected mass () and explosion energy () as a function of progenitor radius (), via the following relations:
where is in units of , , is the plateau luminosity 50 days after explosion and . Here, is measured from the parametric fit to the bolometric lightcurve using the following equation used in Goldberg et al. (2019); Valenti et al. (2016):
where represents the inflection point of the lightcurve and corresponds to the midpoint of the plateau-to-tail transition.
The constraints for LLIIP SNe thus obtained are shown in Figure 11.
6 Results
In this section, we provide a summary of the results derived using the analysis techniques discussed in the previous section.
6.1 Observed Parameters
Parameter | Sample | Median | Range |
Plateau Duration (days) | M | 89 | 64–108 |
M | 75 | 47–137 | |
All | 76 | 47–137 | |
OPTd (days) | M | 99 | 67–116 |
M | 88 | 52–157 | |
All | 90 | 52–157 | |
Rise Time (days) | M | 13 | 10–16 |
M | 13 | 7–22 | |
All | 13 | 7–22 | |
Plateau Slope ( mag day-1) | M | 0.1 | -0.4–0.6 |
M | 0.5 | -0.1–2.0 | |
All | 0.4 | -0.4–2.0 | |
Plateau - Tail Drop (mag) | M | -1.33 | -2.48–0.22 |
M | -1.33 | -3.50—0.09 | |
All | -1.33 | -3.50–0.22 |
The observed parameters in the -band and their errors are shown in Table A. The sample statistics are shown in Table 2. The lower and upper uncertainties correspond to the 16th and 84th percentiles, respectively. The distributions of these observed parameters are shown in Figure 2.
For LLIIP SNe (), the median plateau duration was days, with a range of 64 to 108 days. The median duration of the optically thick phase (OPTd) was days, spanning from 67 to 116 days. The median rise time was days, with a range from 10 to 16 days. The median plateau slope was mag day-1, ranging from to mag day-1. Finally, the change in magnitude between the plateau and the radioactive tail had a median value of mag, with observed values ranging from to mag.
For the full Type IIP SN sample, the median values were days for the plateau duration, days for OPTd, days for the rise time, mag day-1 for the plateau slope, and mag for the plateau-to-tail drop. The steepest drop of mag was seen in SN 2022aagp, shown in Figure 5. A sharp decline from the plateau to the tail is predicted for ECSNe and failed SNe with very low or zero nickel mass (see e.g., Tominaga et al., 2013; Turatto et al., 1998). Other objects with sharp drops include SN 2021tyw ( mag), SN 2022prv ( mag), SN 2023vog ( mag) and SN 2020cxd ( mag).
We further investigated how the lightcurve observables correlate with peak magnitude. As shown in Figure 3, the plateau slope shows a strong correlation with peak brightness, with a Pearson correlation coefficient of (). The best-fit line is given by:
Brighter SNe tend to have steeper plateau slopes, indicating more rapid declines during the plateau phase. We also note that a major fraction of LLIIP SNe have positive slope. The plateau duration shows only a weak correlation with peak magnitude (). The plateau durations for faint and bright SNe span a broad range and overlap significantly. The physical implications for these results are discussed in Section 7.1.
6.2 Progenitor and explosion parameters
6.2.1 Semi-analytical modeling
For the Nagy & Vinkó (2016) models, the nickel mass is only measured for the group of SNe with a distinct radioactive tail. Based on the semi-analytical modeling, the lowest nickel mass was measured for ZTF20aapchqy/SN 2020cxd ( M⊙). Among the LLIIP sample, the median nickel mass is M⊙, with a range of – M⊙, significantly lower than that of the brighter SNe.
The progenitor radii span – R⊙, with a median of R⊙. The ejecta mass ranges from to M⊙, with a median of M⊙. These explosions have a median kinetic energy of erg and values spanning to erg.
For the full Type IIP SN sample, the corresponding median values are: nickel mass M⊙, radius R⊙, ejecta mass M⊙, and kinetic energy erg. The best-fit values for each SN are listed in Table D. The sample statistics are summarized in Table 3.
Parameter | Sample | Median | Range |
() | M | 1.32 | 0.09–4.20 |
M | 7.36 | 0.56–30.98 | |
All | 6.52 | 0.09–30.98 | |
() | M | 8.06 | 5.73–12.89 |
M | 13.90 | 5.41–24.79 | |
All | 12.80 | 5.41–24.79 | |
( erg) | M | 0.79 | 0.21–3.64 |
M | 4.74 | 1.04–9.94 | |
All | 4.32 | 0.21–9.94 | |
() | M | 299 | 138–1024 |
M | 340 | 88–1065 | |
All | 337 | 88–1065 |
6.2.2 Radiation-hydrodynamical modeling
Parameter | Sample | Median | Range |
( erg) | M | 0.17 | 0.10–0.28 |
M | 0.91 | 0.16–4.43 | |
All | 0.76 | 0.10–4.43 | |
( cm) | M | 4.3 | 1.8–8.6 |
M | 5.8 | 1.1–8.4 | |
All | 5.8 | 1.1–8.6 | |
() | M | 1.3 | 0.1–2.5 |
M | 4.1 | 0.2–22.2 | |
All | 3.8 | 0.1–22.2 | |
() | M | 10.5 | 9.4–11.0 |
M | 11.8 | 9.3–16.7 | |
All | 11.5 | 9.3–16.7 | |
) | M | 3.32 | 2.86–4.37 |
M | 3.00 | 1.29–4.48 | |
All | 3.01 | 1.29–4.48 |
For the Moriya et al. (2023) models, the ZAMS mass for the LLIIP SNe is typically M⊙, with a median value of M⊙, and a range of – M⊙. The CSM radii span from to cm, with a median of cm. The explosion energy (Eexp) distribution of LLIIP SNe has a median of erg, and values spanning to erg. The typical pre-supernova mass-loss rates inferred have a median of in units, and a range from to .
For the full sample of Type II SNe, the median ZAMS mass is M⊙, with a CSM radius of cm, explosion energy of erg, and mass-loss rate of in units, as summarized in Table 4.
The best-fit lightcurves are shown in Figure 8. After marginalization, we have estimates with confidence intervals (1) for each of these parameters. The distribution of the best-fit parameters and the correlations are shown in Figure 10 and Figure 14. The best-fit values for each SN are listed in Table E. The sample statistics are summarized in Table 4
6.3 Scaling Relations
We investigate the physical properties of the low-luminosity Type IIP (LLIIP) SNe in our sample using scaling relations between plateau luminosity, nickel mass, duration, and progenitor radius. At a representative radius of R⊙, we find that the LLIIP SNe are characterized by relatively low ejecta masses ranging from to M⊙ and low explosion energies between and erg. These results reinforce the picture that LLIIP SNe arise from low-mass progenitors undergoing weak explosions, consistent with those derived from semi-analytical and radiation-hydrodynamical analysis.
7 Discussion
7.1 Observables
7.1.1 LLIIP SNe exhibit less steep gradients, with most displaying upward slopes
We first attempt to explain the positive slope seen in LLIIP SNe through basic order-of-magnitude arguments. In standard Type IIP models such as Popov (1993), the onset of the plateau is set by hydrogen recombination, where the luminosity reaches
with K being the hydrogen recombination temperature. For low-luminosity plateaus, this implies a smaller ejecta radius at the onset of the plateau phase, since
This is consistent with the lower expansion velocities observed in LLIIP SNe compared to more luminous events, resulting in smaller ejecta radius.
The photon diffusion time at this stage is given by
and can be estimated as
The diffusion time decreases with time as the ejecta expands, owing to both the dilution of density and the recombination-driven reduction in opacity.
Following the explosion, the ejecta cool both adiabatically (on the dynamical timescale ) and radiatively (on the diffusion timescale ). The plateau slope is governed by the ratio of these timescales. When , the internal energy (dominated by radiation) evolves as
As the photosphere recedes, the emergent flux remains approximately blackbody at , with the luminosity given by
Assuming a uniform density profile in the inner ejecta, the mass enclosed within radius evolves as
Combining the above, the photospheric radius evolves as
Using , this simplifies to
and the luminosity becomes
Thus, rising plateaus can naturally emerge in the low-luminosity regime.
The notably flatter or slightly rising slopes in LLIIP SNe can also be explained by the contribution of 56Ni decay to the plateau luminosity. In Figure 12, we show the radial density and mass profiles of – RSG and sAGB progenitors from the literature. The steeper mass density profiles of low-mass progenitors imply that the photospheric radius remains closer to the core of the star throughout the plateau phase (Sato et al., 2024). As a result, 56Ni decay likely makes a stronger contribution to the overall plateau luminosity. Nakar et al. (2016) demonstrated that 56Ni decay can contribute significantly () to the total luminosity during the photospheric phase, influencing the plateau slope. Although their models do not explicitly predict rising slopes, a combination of plateau extension and slope flattening could lead to an apparent increase in brightness over time as seen in SN 2016gfy (Singh et al., 2019). Kozyreva et al. (2019) supports this interpretation, showing that both cooling-envelope emission and nickel distribution shape the late-time behavior of the plateau. However, we note that such 56Ni-driven flattening or rising would not occur unless the progenitor radii were small. For larger progenitors, the recombination-powered luminosity would dominate and naturally decline with time, overpowering the modest nickel contribution. Thus, in the low-luminosity regime, the small radii and low explosion energies of LLIIP progenitors are necessary for 56Ni decay to leave a visible imprint on the plateau slope. Furthermore, Kozyreva et al. (2021) find that ECSNe and low-mass CCSNe from – progenitors exhibit rising light curves in the and redder bands during the first 50 days, followed by a plateau and a subsequent drop to the radioactive tail. This is consistent with the steadily rising -band light curves observed in many LLIIP SNe.
Another possible contributor in the rise of the of the lightcurves could be explosion asymmetry. 3D simulations for a 9 by Stockinger et al. (2020) show that large nickel-rich plumes can distort the ejecta and drive faster expansion in some directions, which lifts the light curve during the initial decline to the plateau and the early phase of the plateau. However, for a large sample, such viewing-angle-dependent effects are expected to average out.
7.1.2 The plateau duration and peak magnitude are not correlated
Single-star models predict a strong correlation between peak brightness and plateau duration in Type II SNe (e.g., Kasen & Woosley, 2009). However, the observed lack of such a correlation suggests that the stars undergo substantial pre-supernova mass loss, likely requiring additional processes such as binary interaction to explain the observed diversity, especially at the low-mass end. Eldridge et al. (2018); Hiramatsu et al. (2021); Dessart et al. (2024) also find that binary stripping can explain the duration of plateau seen in Type II SNe lightcurves. This aligns with the increasing observational consensus that a significant number of massive stars are found in binary systems, with many of them undergoing interaction (e.g., Sana et al., 2012). Similarly, Zapartas et al. (2019) estimate that of all Type II SN progenitors undergo binary interaction. Inferring the explosion and progenitor properties from the lightcurves using models that incorporate self-consistent progenitor evolution with binary effects can be used to constrain the binary fraction and extent of stripping among massive stars that explode as core-collapse SNe.
7.2 Correlation of parameters
The dependence of various physical parameters on peak luminosity is shown in Figures 13 and 14. For the Nagy & Vinkó (2016) semi-analytical models, we find a strong correlation between nickel mass and peak absolute magnitude, with a Pearson coefficient of . This follows a log-linear relation:
Kinetic energy also correlates strongly with peak magnitude (), described by:
The strongest correlation is observed between the ratio of kinetic energy to ejecta mass and the peak magnitude ():
These trends show that lower-luminosity SNe tend to have lower nickel masses, lower explosion energies, and lower energy-to-mass ratios.
The correlations among physical parameters inferred from the Nagy & Vinkó (2016) models are shown in Figure 15. Kinetic energy and nickel mass are strongly correlated (), following:
Kinetic energy also correlates tightly with ejecta mass (), described by:
Nickel mass and ejecta mass show a weaker positive correlation (). Radius is only weakly correlated with peak luminosity (), and pre-SN mass-loss rates appear uncorrelated.
For the Moriya et al. (2023) radiation-hydrodynamical model fits, significant correlations are also observed. Nickel mass correlates with peak magnitude ():
Explosion energy shows a strong correlation ():
The ratio of explosion energy to ejecta mass exhibits the strongest correlation with peak magnitude ():
Explosion energy correlates with nickel mass ():
ZAMS mass correlates moderately with nickel mass (), radius () and Mr (), but with large scatter. Mass-loss rate () shows no significant correlations ().
7.3 Comparison with literature
A direct comparison with previous studies is challenging due to the heterogeneous nature of the Type IIP SN samples analyzed, as well as the differing methodologies used to infer physical parameters. Below, we briefly summarize relevant literature results for key explosion parameters and compare them with values derived from our semi-analytical modeling and radiation-hydrodynamic modeling.
Rodríguez et al. (2021) studied 110 normal Type II SNe, finding 56Ni masses between – M⊙, with an average of M⊙. Martinez et al. (2022a) reported 56Ni masses of – M⊙ (median M⊙) from 17 objects. Pejcha & Prieto (2015) and Müller et al. (2017) identified similar medians ( M⊙), though their maximum values reached up to M⊙. Meanwhile, Subrayan et al. (2023) measured the nickel mass for 45 SNe, with the lowest nickel mass for one SN being 0.01 and the others ranging between 0.02 and 0.1 . Anderson (2019) measured a broader range of – M⊙, also with a median near M⊙. Hamuy (2003) measured 0.0016 to 0.26 for a sample of 21 Type II SNe. Our semi-analytical modeling yields a high median nickel mass of for the full Type IIP SN sample, spanning 0.0009–0.3098 . The LLIIP subsample shows significantly lower nickel masses, with a median of and a range of 0.0009–0.0420 . The radiation-hydrodynamical fits (Moriya et al., 2023) yield broadly consistent results, with a median nickel mass of for the full sample and for the LLIIP subsample. These results imply that LLIIP SNe are characterized by relatively lower 56Ni yields, in agreement with earlier studies and theoretical predictions for the explosion of low-mass cores (see Section 7.5).
Observationally derived explosion energies typically cluster around – erg (e.g., Martinez et al., 2022a). Martinez et al. (2022a) specifically report a median explosion energy of erg (range – erg). Hamuy (2003) measured erg for a sample of 21 Type II SNe. Utrobin & Chugai (2019) and Pumo et al. (2017) report higher energies, up to – erg. Our semi-analytical modeling yields a high median explosion energy of erg for the full sample, with a range from 0.21–9.94 erg. For the LLIIP subsample, the median energy is substantially lower at erg. We note that the ejecta masses and explosion energies from the semi-analytical fits from Nagy & Vinkó (2016) are higher than those measured in the literature. This is likely due to the two-component density profile assumed in these models. The caveats of the models are further discussed in Section 7.4. The radiation-hydrodynamical fits (Moriya et al., 2023) show relatively lower energies. For the full sample, the median energy is erg, and for the LLIIP events, it is erg, ranging from 0.10–0.28 erg. We note that the lowest energy in the model grid was erg, so it is possible that the explosion energies are even lower.
Typical ejecta masses reported by previous studies, such as Martinez et al. (2022a), are around M⊙ (16–84 percentile range is M⊙). Utrobin & Chugai (2019) report higher ejecta masses, up to M⊙. Hamuy (2003) reported even higher ejecta masses, ranging from . ZAMS masses measured in Subrayan et al. (2023) and Silva-Farfán et al. (2024) ranged from , with most objects at M☉ and a very steep IMF. Our semi-analytical modeling yields a median ejecta mass of for the full sample, with a range from 5.41–24.79 . The LLIIP subsample shows a lower median ejecta mass of , still broadly consistent with the lower end of the literature estimates. This also suggests that the hydrogen envelopes of these massive stars are stripped before they undergo core-collapse. The radiation-hydrodynamical fits (Moriya et al., 2023) give a median ZAMS mass of for the full sample, and for the LLIIP SNe.
Typical progenitor radii reported by previous studies are around – R⊙, with a median of R⊙ (Martinez et al., 2022a). Progenitor radii of 80 to 600 was measured in Hamuy (2003). Irani et al. (2024) find relatively higher radius, with 15 out of 34 SNe II well fit to models with breakout radii cm, while 18 SNe are better fit by a model with a large cm breakout radius. Our semi-analytical modeling yields somewhat smaller radii, with a median of for the full sample and for the LLIIP subsample. These results suggest that LLIIP progenitors may originate from more compact red supergiants or that their outer envelopes have been stripped off via binary interaction at the onset of core-collapse. For comparison, the radius distribution of 74 Galactic RSGs spans a broad range from to R⊙, with a median of R⊙ (Levesque et al., 2005).
Our radiation-hydrodynamical modeling yields median mass-loss rates of for the full sample and for the LLIIP subsample. LLIIP SNe progenitors likely have lower mass-loss prior to explosion. Our results are consistent with the mass-loss rates inferred by Subrayan et al. (2023) and Silva-Farfán et al. (2024), who report values in the range – . The inferred CSM radii have a median value of cm for the full sample and cm for LLIIP events, indicating extended circumstellar environment that could influence early-time lightcurve behavior (e.g., Galbany et al., 2016; Valenti et al., 2016; Förster et al., 2018; Morozova et al., 2018; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2024; Hinds et al., 2025; Jacobson-Galán et al., 2025).
7.4 Caveats of the models
We caution that while the models employed to fit the lightcurves provide useful constraints and correlations, they also introduce several assumptions and systematic effects that may influence the derived parameters. We summarize these below.
A key challenge arises from the degeneracy between key parameters such as ejecta mass, explosion energy, and progenitor radius. Fitting models to photometry in only a couple of passbands allows multiple combinations of physical parameters to produce similar light curves. This degeneracy limits our ability to extract unique solutions from lightcurve fitting alone as discussed in Singh et al. (2024); Rehemtulla et al. (2025). Goldberg et al. (2019); Goldberg & Bildsten (2020) argue that breaking these degeneracies requires external constraints on at least one physical parameter.
Moriya et al. (2023) attempt to mitigate this degeneracy using the progenitor mass–radius relation embedded in the KEPLER stellar grids. Although their lightcurve model grid spans a wide range of progenitor and explosion properties, it is important to note that there are systematic limitations inherited from the underlying stellar evolution framework. For example, they employ a low mixing-length parameter, leading to progenitor radii that are systematically larger than what is inferred from observational constraints. Additionally, the assumption of continuous wind-driven mass loss throughout stellar evolution results in hydrogen-rich envelope masses that are systematically lower at the time of core collapse. These can influence the lightcurves and the physical parameters derived from fitting them (Hsu et al., 2024). While we incorporate velocity information from photospheric-phase spectroscopy to partially constrain kinetic energy, these measurements are not as constraining as those obtained during the brief shock-cooling phase immediately after explosion, which are more diagnostic (Goldberg et al., 2019). We also note that the lowest explosion energy in the model grid is erg; lower energies are required to better fit the lightcurves of the faintest SNe in our sample.
The semi-analytical two-component lightcurve models developed by Nagy & Vinkó (2016) assume spherical symmetry and do not accurately capture the early shock-cooling behavior at days. Accordingly, we restrict our fits to data obtained after 20 days post-explosion. Despite their simplicity, Nagy & Vinkó (2016) and Jäger et al. (2020) find that these models yield physical parameters broadly consistent with more sophisticated hydrodynamical simulations for Type IIP SNe. To test this consistency, Yang et al. (2021) compared fitting results to those from hydrodynamical models such as those in Martinez et al. (2020), and found that, in several cases, the inferred ejecta masses are in reasonable agreement. We find that the semi-analytical fits return higher explosion energies and ejecta mass for most SNe. This could be because of the two-component density profile and a simplified method for computing bolometric lightcurves. Prior works have also noted that lightcurve modeling can systematically overestimate ZAMS masses compared to direct progenitor imaging or nebular spectroscopy (e.g., Utrobin & Chugai, 2008, 2009; Maguire et al., 2010; Sanders et al., 2015).
The correlations between physical quantities such as peak luminosity, nickel mass, explosion energy, and ejecta mass are preserved in both approaches. Nevertheless, we also observe that, for individual SN, the absolute values derived from the two methods can differ significantly (Figure 21). This highlights the presence of methodological systematics and the need for caution when interpreting parameter estimates from any single modeling framework.
7.5 What are the progenitors of LLIIP SNe?
7.5.1 7–12 Red Super Giants that undergo core-collapse SNe
We compare our derived explosion parameters to theoretical predictions from neutrino-driven explosion models of core-collapse SNe. Barker et al. (2022) show that the properties of a progenitor’s core can be estimated from optical photometry during the plateau phase alone. A linear correlation between iron core mass and (luminosity 50 days post-explosion) indicates that optical photometry of Type IIP SNe can effectively probe progenitor cores. They find a robust relationship linking progenitor iron core mass to plateau luminosity (see Figure 17). The bolometric luminosities of our LLIIP SN sample align with ZAMS masses below 12 . However, we note that these models assume that there is no CSM interaction. Additionally, using best-fit ZAMS mass estimates from Moriya et al. (2023) mode fits, the ZAMS mass for LLIIP SNe are lower than 11 . Furthermore, ejecta masses inferred from semi-analytical model fitting by Nagy & Vinkó (2016) suggest masses 8.2 . Assuming a neutron star mass of 1.4 , this translates to progenitor masses less than .
Also, Recent 3D simulations by Burrows et al. (2020) find explosion energies of approximately 0.1 foe for their lowest ZAMS mass (9 ) progenitor model, compatible with our lowest observed explosion energies from the radiation-hydrodynamical fits (see Figure 17). Previous 1D simulations (e.g., Ugliano et al., 2012; Ertl et al., 2016; Sukhbold et al., 2016) yielded explosion energies between 0.1 and 2.0 foe.
Observations and parameterized 1D explosion models suggest typical nickel yields of from core-collapse SNe. However, the low nickel masses we measure imply progenitor initial masses toward the low-mass end of the CCSN spectrum. Stars with low-mass iron cores are known to produce significantly reduced nickel yields. For example, Stockinger et al. (2020) and Sandoval et al. (2021) find nickel masses of 0.002–0.005 for a 9.6 progenitor. Recent 3D simulations by Burrows et al. (2024) predict nickel masses in the range 0.002–0.006 for a 9 progenitor, rising above 0.01 for progenitor masses exceeding 9.25 . These theoretical predictions are consistent with our measured nickel mass for LLIIP SNe.
7.5.2 sAGB stars that undergo Electron-capture SNe
Progenitors in the mass range may evolve into super-Asymptotic Giant Branch (sAGB) stars and explode as electron-capture supernovae (ECSNe). ECSN simulations consistently produce low explosion energies and nickel yields significantly lower than typical Fe CCSNe but similar to those observed for the lowest-mass Fe CCSNe. According to Sato et al. (2024), ECSNe, without significant circumstellar material (CSM) interaction, exhibit distinctively bluer bolometric lightcurve plateaus compared to Fe CCSNe, characterized by mid-plateau colors exceeding the relation . This difference arises from the lower-density envelopes inherent to sAGB progenitors. We plot the mid-plateau colors for our SN sample against the Sato et al. (2024) criterion in Figure 18. While SNe 2022omr, 2021tyw, 2021cwe, and 2022prv show bluer mid-plateau colors, their peak luminosities and nickel masses favor an Fe CCSN origin. Among our LLIIP SNe sample, SNe 2023wcr, 2021gmj, 2021zgm, and 2020abcq exhibit notably bluer mid-plateau colors. However, we note that the mid-plateau colors might also be affected by significant CSM interaction. The lightcurves of the SNe with the bluest (SN 2021tyw, mid-plateau mag) and reddest (SN 2022jzc, mid-plateau mag) colors are shown in Figure 18.
We also compare the inferred nickel masses and explosion energies of our LLIIP sample to the ECSN models from Kozyreva et al. (2021). The nickel masses () and explosion energies () from our sample closely match those predicted by ECSN simulations for their 8.8 and 9.6 progenitor models. Kozyreva et al. (2021) also show that the lightcurves in the and bands rises during the first 50 days and then slowly declines, while the light curve in the and redder bands rises over the same timescale and then settles into a plateau before a steep drop to the radioactive tail. The rising -band light curves seen in some LLIIP SNe, when coupled with their low nickel masses, are consistent with theoretical expectations for explosions from sAGB stars that undergo ECSNe.
Photospheric properties alone do not reliably distinguish between ECSNe from sAGB stars and Fe-core collapse SNe from red supergiants. Nebular-phase spectroscopy, however, offers definitive diagnostic features due to distinct core compositions and nucleosynthetic pathways (Jerkstrand et al., 2018). A comprehensive analysis of nebular spectra for the LLIIP SN sample will be presented in Paper III.
8 Conclusion
We present the largest systematic study of lightcurve properties for Type IIP SNe from the volume-limited ZTF CLU survey, analyzing 129 Type IIP events, including 16 LLIIP SNe. We summarize the key takeaways of the paper here:
1.
Observables:
The median plateau duration for LLIIP SNe is days, compared to days for the full Type IIP sample. The median OPTd is days for LLIIP SNe and days for the full sample. There is no strong correlation with peak magnitude, likely suggesting binary interaction.
The plateau slope strongly correlates with peak brightness (), with brighter SNe having steeper declines; a significant fraction of LLIIP SNe exhibit positive plateau slopes. This likely implies progenitors for most LLIIP SNe have a smaller radius and different density profile from those of their more luminous counterparts.
We observe a strong correlation between the mid-plateau luminosity and nickel mass (), extending to the low-luminosity end. This can be explained if a strong core explosion leads to both higher explosion energy and greater nickel yield.
The steepest plateau-to-tail drop of mag was seen in SN 2022aagp. Other objects with sharp drops include SN 2021tyw ( mag) and SN 2020cxd ( mag). These are consistent with ECSNe and failed SNe with very low or zero nickel mass.
2.
Explosion and Progenitor Analysis:
Radiation-hydrodynamical modeling: The ZAMS masses for LLIIP SNe are typically below M⊙, with a median of M⊙, while that of the overall sample is M⊙. The kinetic energies are similarly lower for LLIIP SNe, with a median of erg, compared to erg for the overall sample. The inferred mass-loss rates just before explosion are relatively low, with a median for LLIIP SNe, compared to for the overall sample.
Semi-analytical modeling: LLIIP SNe have low nickel masses, with a median of M⊙ and a range from to M⊙, while that of the overall Type IIP SNe sample is M⊙ with a range from to M⊙. The median ejecta mass for LLIIP SNe is M⊙ compared to M⊙ for the overall sample. Their kinetic energies are also lower, with a median of erg for LLIIP SNe and erg for the overall sample.
Overall, LLIIP SNe occupy the faint, low-energy, low-mass end of the Type IIP population compared to more luminous events.
3.
Correlation of Parameters:
Nickel mass, kinetic and explosion energy, and energy-to-mass ratio strongly correlate with peak absolute magnitude, indicating that fainter SNe have systematically lower nickel mass and explosion energies. This is consistent with theoretical core-collapse SN models, which predict a strong correlation between explosion energy and nickel production.
For the semi-analytical fits, nickel mass and kinetic energy correlates with peak brightness ( and respectively). Kinetic energy-to-ejecta mass ratio shows the strongest correlation () with peak brightness. Among physical parameters themselves, kinetic energy correlates strongly with ejecta mass () and with nickel mass ().
For the radiation-hydrodynamical fits, similar trends are observed: peak brightness correlates tightly with explosion energy (), nickel mass () and energy-to-ejecta mass ratio (). Explosion energy also correlates with nickel mass (), while mass-loss rate shows little dependence on other explosion parameters.
The correlations between physical parameters such as peak luminosity, nickel mass, explosion energy, and ejecta mass are preserved across both semi-analytical and radiation-hydrodynamic modeling approaches. Nevertheless, we observe that for individual SNe, the absolute values derived from the two methods can differ significantly. This highlights the presence of methodological systematics and underscores the need for caution when interpreting parameter estimates from any single modeling framework. However, the low explosion energies, nickel masses and progenitor masses inferred for LLIIP SNe are consistent with the explosions of low-mass RSG or sAGB stars with ZAMS masses 12 M⊙. Distinguishing between these progenitor scenarios will require late-time nebular spectroscopy, which will be pursued in Paper III of this series. The models in this paper assumes single-star evolution. Incorporating progenitor models that account for binary interaction and mass transfer will be crucial for constraining the binary fraction and understanding the diversity of envelope stripping among core-collapse progenitors. We also highlight the need for radiation-hydrodynamic model grids that extend to the lowest progenitor masses in the 8–10 M⊙ range and explore explosion energies below erg to adequately model the faintest Type IIP SNe. The advent of deep, multi-band, high-cadence surveys such as the Legacy Survey of Space and Time (LSST; Ivezić et al., 2008) will enable a more complete census of the faintest slow-evolving core-collapse SNe and provide critical insights into the lowest-mass stars that explode. It will also offer deep post-plateau limits, facilitating the discovery of more Type IIP SNe with very low or even zero nickel mass, as expected for electron-capture supernovae or failed explosions.
Data Availability
All photometric and spectroscopic data presented in this work will be made publicly available on Zenodo and WISeREP upon publication. The codes required to reproduce the lightcurve fitting will be uploaded via GitHub. Machine-readable versions of all tables, lightcurve plots, and best-fit model plots will also be available on Zenodo.
9 Acknowldegement
We thank Jared Goldberg, Daichi Hiramatsu and Kishalay De for valuable discussions.
Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341, AST-2034437, and currently Award 2407588. ZTF receives additional funding from the ZTF partnership. Current members include Caltech, USA; Caltech/IPAC, USA; University of Maryland, USA; University of California, Berkeley, USA; University of Wisconsin at Milwaukee, USA; Cornell University, USA; Drexel University, USA; University of North Carolina at Chapel Hill, USA; Institute of Science and Technology, Austria; National Central University, Taiwan, and OKC, University of Stockholm, Sweden. Operations are conducted by Caltech’s Optical Observatory (COO), Caltech/IPAC, and the University of Washington at Seattle, USA.
Zwicky Transient Facility access for S.S. was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA).
M.W.C acknowledges support from the National Science Foundation with grant numbers PHY-2117997, PHY-2308862 and PHY-2409481.
D.T. is supported by the Sherman Fairchild Postdoctoral Fellowship at Caltech.
E.O. is supported by the Swedish Research Council (Project No. 2020-00452).
N.Sarin acknowledges support from the Knut and Alice Wallenberg Foundation through the “Gravity Meets Light” project and by and by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012.
SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171.
The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant #12540303 (PI: Graham).
The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal .
This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council.
The W. M. Keck Observatory is operated as a scientific partnership among the California Institute of Technology, the University of California and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
Software: Global Relay of Observatories Watching Transients Happen Marshal (GROWTH; Kasliwal et al., 2019) and the Fritz SkyPortal Marshal (Duev et al., 2019; van der Walt et al., 2019; Coughlin et al., 2023). Astropy (Astropy Collaboration et al., 2013), Matplotlib (Hunter, 2007), george (Ambikasaran et al., 2015)
Appendix
Table A lists the observable lightcurve parameters such as rise time, plateau duration, and decline rate for the entire sample. Tables B and C summarize the priors used in the semi-analytical and radiation-hydrodynamical model fits, respectively. Tables D and E present the full posterior estimates of explosion and progenitor properties from the two modeling approaches. Table F provides the spectral observation log and H velocity measurements. Figure 19 shows all the parameter correlations based on semi-analytical fits, while Figure 20 shows the same using radiation-hydrodynamical models. Figure 21 compares results between the two modeling methods.
ZTF | IAU | Rise | Plat. dur. | OPTd | Slope | mag | Limit? | |
(mag) | (days) | (days) | (days) | (0.01 mag day-1) | (mag) | |||
ZTF24abtczty | SN 2024abfl | -14.1 | 12.2 ± 0.4 | 99.1 ± 1.5 | 111.7 ± 1.5 | -0.04 ± 0.02 | -2.05 | no |
ZTF20aapchqy | SN 2020cxd | -14.2 | 108.1 ± 1.6 | 112.9 ± 1.6 | -0.42 ± 0.02 | -2.48 | no | |
ZTF24abmkros | SN 2024xkd | -14.9 | 90.5 ± 1.3 | 97.1 ± 1.3 | -0.11 ± 0.02 | -1.17 | yes | |
ZTF23aackjhs | SN 1995al | -15.0 | 97.0 ± 2.3 | 99.5 ± 2.3 | 0.14 ± 0.02 | 0.22 | yes | |
ZTF22aakdbia | SN 2022jzc | -15.1 | 9.7 ± 0.3 | 92.8 ± 5.5 | 111.0 ± 5.5 | -0.18 ± 0.02 | -2.28 | no |
ZTF22abtjefa | SN 2022aaad | -15.1 | 16.0 ± 0.3 | 66.5 ± 15.0 | 85.8 ± 15.0 | -0.02 ± 0.03 | -1.38 | no |
ZTF22aazmrpx | SN 2022raj | -15.2 | 80.0 ± 4.2 | 84.7 ± 4.2 | 0.45 ± 0.03 | -0.87 | yes | |
ZTF20acuhren | SN 2020abcq | -15.3 | 101.2 ± 2.8 | 103.2 ± 2.8 | -0.01 ± 0.02 | -2.05 | yes | |
ZTF21aaobkmg | SN 2021eui | -15.4 | 85.0 ± 3.1 | 88.0 ± 3.1 | 0.09 ± 0.02 | -1.43 | yes | |
ZTF22aaywnyg | SN 2022pru | -15.4 | 86.1 ± 1.6 | 102.1 ± 1.6 | 0.46 ± 0.02 | -1.09 | no | |
ZTF24aabppgn | SN 2024wp | -15.5 | 95.3 ± 2.5 | 109.4 ± 2.4 | 0.1 ± 0.02 | -1.35 | yes | |
ZTF23abnogui | SN 2023wcr | -15.6 | 87.3 ± 4.3 | 101.1 ± 4.3 | 0.58 ± 0.04 | -1.53 | no | |
ZTF24aaucrua | SN 2024nez | -15.8 | 90.2 ± 1.5 | 105.4 ± 1.5 | 0.18 ± 0.02 | -1.58 | yes | |
ZTF22abyivoq | SN 2022acko | -15.8 | 69.8 ± 6.4 | 84.7 ± 6.4 | 0.23 ± 0.04 | 0.02 | yes | |
ZTF20abeohfn | SN 2020mjm | -15.9 | 10.7 ± 0.1 | 64.3 ± 7.6 | 76.5 ± 7.6 | 0.05 ± 0.03 | 0.05 | yes |
ZTF22abssiet | SN 2022zmb | -15.9 | 66.1 ± 11.3 | 76.0 ± 11.3 | 0.18 ± 0.04 | -0.24 | yes | |
ZTF19aamwhat | SN 2019bzd | -16.0 | 83.1 ± 1.5 | 103.1 ± 1.4 | 0.36 ± 0.02 | -2.00 | yes | |
ZTF22abkhrkd | SN 2022wol | -16.0 | 79.1 ± 2.6 | 90.8 ± 2.6 | 0.08 ± 0.03 | -1.09 | yes | |
ZTF24aaejecr | SN 2024btj | -16.0 | 78.3 ± 2.3 | 89.6 ± 2.3 | 0.55 ± 0.03 | -1.66 | no | |
ZTF24aaplfjd | SN 2024jxm | -16.0 | 74.6 ± 4.6 | 97.6 ± 4.6 | 0.24 ± 0.03 | -1.30 | no | |
ZTF19abwztsb | SN 2019pjs | -16.2 | 68.1 ± 37.2 | 80.0 ± 37.2 | 0.37 ± 0.2 | -1.36 | no | |
ZTF24aabsmvc | SN 2024ws | -16.2 | 90.0 ± 23.0 | 100.4 ± 23.0 | 0.26 ± 0.07 | -1.25 | yes | |
ZTF22abyohff | SN 2022acrl | -16.3 | 78.7 ± 6.5 | 88.7 ± 6.4 | 0.33 ± 0.04 | -0.52 | yes | |
ZTF21acgrrnl | SN 2021aayf | -16.3 | 77.0 ± 14.8 | 90.5 ± 14.8 | -0.14 ± 0.04 | -1.44 | yes | |
ZTF21aaeqwov | AT 2021htp | -16.4 | 13.2 ± 0.6 | 92.4 ± 1.1 | 106.1 ± 1.1 | 0.27 ± 0.02 | -1.30 | no |
ZTF23abaxtlq | SN 2023rix | -16.4 | 75.2 ± 8.3 | 87.3 ± 8.3 | 0.92 ± 0.1 | -2.13 | no | |
ZTF21abvcxel | SN 2021wvw | -16.4 | 67.4 ± 1.2 | 69.4 ± 1.2 | 0.47 ± 0.03 | -1.67 | yes | |
ZTF21aantsla | SN 2021ech | -16.4 | 75.3 ± 8.0 | 86.5 ± 8.0 | 0.25 ± 0.04 | -1.46 | yes | |
ZTF21aafepon | SN 2021ass | -16.4 | 51.4 ± 9.0 | 67.9 ± 9.0 | 0.16 ± 0.05 | -0.09 | yes | |
ZTF21acpqqgu | SN 2021aewn | -16.4 | 76.3 ± 9.7 | 78.3 ± 9.7 | 0.35 ± 0.05 | -0.77 | yes | |
ZTF22abnujbv | SN 2022xus | -16.4 | 53.2 ± 20.7 | 62.4 ± 20.7 | 0.58 ± 0.23 | -1.83 | no | |
ZTF20acmaaan | SN 2020xyk | -16.4 | 69.4 ± 4.2 | 75.8 ± 4.2 | 0.14 ± 0.03 | -1.96 | yes | |
ZTF22abfxkdm | SN 2022ubb | -16.6 | 87.3 ± 3.9 | 96.2 ± 3.9 | 0.49 ± 0.03 | -1.66 | yes | |
ZTF22aavbfhz | AT 2022phi | -16.6 | 78.3 ± 2.0 | 81.3 ± 2.0 | 0.63 ± 0.03 | -1.02 | yes | |
ZTF22abkbjsb | SN 2022vym | -16.7 | 60.0 ± 9.5 | 66.5 ± 9.5 | 0.62 ± 0.1 | -1.33 | no | |
ZTF24aabpzuz | SN 2024vs | -16.7 | 62.0 ± 4.7 | 78.9 ± 4.7 | -0.13 ± 0.03 | -1.33 | yes | |
ZTF21aanzcuj | SN 2021enz | -16.7 | 18.0 ± 0.8 | 72.1 ± 1.9 | 90.6 ± 1.8 | 0.01 ± 0.03 | -1.54 | no |
ZTF24aafqzur | SN 2024daa | -16.7 | 67.6 ± 1.2 | 86.6 ± 1.2 | 0.15 ± 0.03 | -0.94 | yes | |
ZTF21aagtqna | SN 2021brb | -16.8 | 21.2 ± 1.4 | 73.6 ± 3.7 | 104.8 ± 3.7 | 0.12 ± 0.03 | -1.26 | no |
ZTF18aaszvfn | SN 2021iaw | -16.8 | 76.8 ± 1.7 | 86.4 ± 1.7 | 0.39 ± 0.03 | -1.23 | yes | |
ZTF19acftfav | SN 2019ssi | -16.8 | 14.2 ± 0.4 | 61.7 ± 8.0 | 76.4 ± 8.0 | 0.71 ± 0.1 | -1.70 | yes |
ZTF23aasbvab | SN 2023ngy | -16.8 | 92.7 ± 1.1 | 102.9 ± 1.1 | 0.83 ± 0.02 | -1.32 | yes | |
ZTF23aasrcyv | SN 2023nlu | -16.8 | 18.3 ± 1.2 | 71.8 ± 1.8 | 94.0 ± 1.8 | 0.5 ± 0.03 | -0.24 | yes |
ZTF22abzqwmp | SN 2022adth | -16.8 | 61.0 ± 11.1 | 73.6 ± 11.1 | 0.69 ± 0.13 | -0.81 | yes | |
ZTF22aasojye | SN 2022omr | -16.9 | 98.1 ± 4.2 | 103.6 ± 4.2 | 0.23 ± 0.02 | -1.76 | no |
ZTF | IAU | Rise | Plat. dur. | OPTd | Slope | mag | Limit? | |
(mag) | (days) | (days) | (days) | (0.01 mag day-1) | (mag) | |||
ZTF19aadnxnl | SN 2019va | -16.9 | 103.2 ± 3.8 | 107.2 ± 3.8 | -0.09 ± 0.02 | -0.80 | no | |
ZTF19actnyae | SN 2019vdm | -16.9 | 52.9 ± 1.4 | 61.4 ± 1.4 | -0.02 ± 0.04 | -1.33 | no | |
ZTF23aaxadel | SN 2023pbg | -16.9 | 89.8 ± 6.1 | 104.6 ± 6.1 | 0.31 ± 0.03 | -0.82 | yes | |
ZTF21abrluay | SN 2021vfh | -17.0 | 75.4 ± 8.8 | 88.0 ± 8.8 | 0.08 ± 0.03 | -1.28 | yes | |
ZTF22aafsqud | SN 2022hql | -17.0 | 61.2 ± 1.6 | 71.7 ± 1.6 | 0.62 ± 0.04 | -1.25 | yes | |
ZTF22abhsxph | SN 2022vyc | -17.0 | 80.3 ± 1.4 | 94.6 ± 1.4 | 0.51 ± 0.03 | -1.42 | yes | |
ZTF19aazudta | SN 2019hqm | -17.0 | 73.7 ± 16.5 | 90.0 ± 16.5 | 0.18 ± 0.05 | -1.52 | yes | |
ZTF22abfwxtr | SN 2022udq | -17.0 | 99.3 ± 2.3 | 111.4 ± 2.3 | 0.41 ± 0.02 | -0.73 | yes | |
ZTF21aafkwtk | SN 2021apg | -17.0 | 69.3 ± 6.0 | 87.0 ± 6.0 | 0.33 ± 0.04 | -1.65 | no | |
ZTF19aapafit | SN 2019cvz | -17.0 | 82.8 ± 2.2 | 98.3 ± 2.2 | 0.12 ± 0.02 | -1.30 | no | |
ZTF23abbtkrv | SN 2023rvo | -17.0 | 16.7 ± 1.1 | 90.7 ± 2.2 | 157.1 ± 2.2 | -0.04 ± 0.02 | -0.81 | yes |
ZTF19acewuwn | SN 2019ssl | -17.0 | 62.8 ± 8.0 | 83.8 ± 8.0 | 0.19 ± 0.04 | -1.09 | yes | |
ZTF19aailepg | SN 2019amt | -17.1 | 100.9 ± 10.3 | 109.0 ± 10.3 | 0.3 ± 0.04 | -1.04 | yes | |
ZTF19aazyvub | SN 2019hnl | -17.1 | 74.6 ± 5.1 | 80.6 ± 5.0 | 0.39 ± 0.04 | -1.44 | no | |
ZTF21aakvroo | SN 2021cwe | -17.1 | 72.5 ± 4.2 | 76.5 ± 4.2 | -0.08 ± 0.03 | -0.93 | no | |
ZTF22abtspsw | SN 2022aagp | -17.1 | 9.0 ± 0.0 | 111.5 ± 6.3 | 122.5 ± 6.3 | 0.99 ± 0.06 | -3.50 | yes |
ZTF18aaxkqgy | SN 2018ccb | -17.1 | 70.0 ± 10.6 | 74.2 ± 10.6 | 0.53 ± 0.09 | -1.22 | no | |
ZTF22abfavpu | SN 2022tmb | -17.1 | 66.6 ± 1.3 | 80.9 ± 1.3 | 0.77 ± 0.03 | -2.05 | no | |
ZTF21aapliyn | SN 2021foj | -17.1 | 91.3 ± 1.8 | 96.4 ± 1.8 | 0.15 ± 0.02 | -1.20 | yes | |
ZTF21acafqtj | SN 2021yok | -17.1 | 59.9 ± 10.6 | 69.9 ± 10.6 | 0.37 ± 0.07 | -0.20 | yes | |
ZTF21abfiuqf | SN 2021pla | -17.2 | 52.9 ± 1.2 | 60.6 ± 1.2 | 1.25 ± 0.05 | -1.51 | no | |
ZTF19aanhhal | SN 2019cec | -17.2 | 80.9 ± 2.8 | 90.2 ± 2.8 | 0.44 ± 0.03 | -0.96 | yes | |
ZTF22abyokkf | SN 2022acri | -17.2 | 91.7 ± 5.2 | 108.2 ± 5.2 | 0.08 ± 0.02 | -1.14 | yes | |
ZTF18aatyqds | SN 2018btl | -17.2 | 10.8 ± 0.2 | 76.5 ± 4.4 | 88.8 ± 4.4 | 1.2 ± 0.07 | -1.39 | no |
ZTF22aapargp | SN 2022niw | -17.2 | 54.0 ± 9.6 | 68.7 ± 9.6 | 0.46 ± 0.09 | -1.03 | yes | |
ZTF23aanymcl | SN 2023kzz | -17.2 | 60.6 ± 2.1 | 64.1 ± 2.1 | 0.68 ± 0.04 | -1.23 | yes | |
ZTF21aavhnpk | SN 2021jsf | -17.3 | 74.3 ± 9.1 | 78.8 ± 9.0 | 0.7 ± 0.09 | -1.34 | yes | |
ZTF19aaniore | SN 2019ceg | -17.3 | 87.2 ± 3.0 | 102.0 ± 3.0 | 0.44 ± 0.03 | -1.46 | no | |
ZTF21aabygea | SN 2021os | -17.3 | 14.8 ± 0.4 | 83.3 ± 1.8 | 100.1 ± 1.8 | 0.58 ± 0.03 | -1.43 | no |
ZTF22absqhkw | SN 2022zkc | -17.4 | 50.1 ± 4.6 | 52.1 ± 4.6 | 0.62 ± 0.07 | -1.48 | yes | |
ZTF23aaqknaw | SN 2023lzn | -17.4 | 100.7 ± 1.7 | 104.7 ± 1.6 | 0.55 ± 0.02 | -2.03 | yes | |
ZTF21aaagypx | SN 2021V | -17.4 | 137.2 ± 1.1 | 146.1 ± 1.1 | 0.56 ± 0.02 | -1.57 | yes | |
ZTF19abbnamr | SN 2019iex | -17.4 | 94.1 ± 10.0 | 100.6 ± 10.0 | 0.17 ± 0.03 | -1.58 | no | |
ZTF19aawgxdn | SN 2019gmh | -17.4 | 95.9 ± 8.0 | 105.5 ± 8.0 | 0.34 ± 0.04 | -1.54 | yes | |
ZTF24aagupsf | SN 2024egd | -17.5 | 71.0 ± 7.0 | 82.4 ± 7.0 | 0.62 ± 0.07 | -2.46 | no | |
ZTF18abzrgim | SN 2018gvt | -17.5 | 73.0 ± 24.1 | 90.4 ± 24.1 | -0.03 ± 0.03 | -0.31 | yes | |
ZTF19acytcsg | SN 2019wvz | -17.6 | 77.4 ± 9.2 | 93.1 ± 9.2 | 0.25 ± 0.04 | -1.13 | no | |
ZTF18aawpwlf | SN 2020hvn | -17.6 | 54.1 ± 3.7 | 62.1 ± 3.7 | 1.76 ± 0.13 | -1.49 | yes | |
ZTF23abhzfww | SN 2023twg | -17.6 | 90.0 ± 3.0 | 96.0 ± 3.0 | 0.6 ± 0.03 | -1.45 | no | |
ZTF24aarvbxj | SN 2024lby | -17.6 | 49.1 ± 1.4 | 60.4 ± 1.4 | 0.73 ± 0.05 | -1.23 | no | |
ZTF19aclobbu | SN 2019twk | -17.6 | 73.7 ± 10.3 | 84.5 ± 10.3 | 0.31 ± 0.05 | -0.52 | yes | |
ZTF19actnwtn | SN 2019vdl | -17.6 | 84.1 ± 1.7 | 102.1 ± 1.7 | 0.39 ± 0.03 | -1.09 | yes | |
ZTF21abouuat | SN 2021ucg | -17.6 | 81.3 ± 11.4 | 98.7 ± 11.4 | 0.27 ± 0.05 | -1.20 | yes | |
ZTF24aaabbse | SN 2023achj | -17.8 | 103.9 ± 9.5 | 123.8 ± 9.5 | 0.65 ± 0.06 | -1.03 | yes | |
ZTF21acgunkr | SN 2021aaxs | -17.8 | 77.7 ± 16.3 | 89.5 ± 16.3 | 0.99 ± 0.21 | -0.52 | yes |
ZTF | IAU | Rise | Plat. dur. | OPTd | Slope | mag | Limit? | |
(mag) | (days) | (days) | (days) | (0.01 mag day-1) | (mag) | |||
ZTF21abnlhxs | SN 2021tyw | -17.8 | 8.7 ± 0.2 | 104.5 ± 1.1 | 116.1 ± 1.1 | 0.85 ± 0.02 | -3.10 | yes |
ZTF19abajxet | SN 2019hyk | -17.8 | 10.9 ± 0.2 | 67.3 ± 10.9 | 79.7 ± 10.9 | 0.95 ± 0.16 | -0.88 | yes |
ZTF19acbwejj | SN 2019upq | -17.8 | 22.0 ± 7.4 | 87.7 ± 3.3 | 130.3 ± 3.3 | 1.14 ± 0.05 | -1.28 | no |
ZTF23abmoxlu | SN 2023vog | -17.8 | 10.9 ± 0.2 | 109.4 ± 2.4 | 123.2 ± 2.4 | 1.19 ± 0.03 | -2.49 | no |
ZTF21abhhrpj | SN 2021qiu | -17.9 | 90.7 ± 1.6 | 96.3 ± 1.6 | 0.53 ± 0.02 | -1.93 | no | |
ZTF22aakdqqg | SN 2022kad | -17.9 | 15.5 ± 0.6 | 63.1 ± 12.8 | 79.6 ± 12.8 | 1.38 ± 0.28 | -1.27 | no |
ZTF23aailjjs | SN 2023hcp | -17.9 | 60.2 ± 2.5 | 69.6 ± 2.4 | 0.83 ± 0.05 | -1.57 | no | |
ZTF19aanrrqu | SN 2019clp | -18.0 | 49.6 ± 4.5 | 53.8 ± 4.5 | 1.17 ± 0.11 | -1.65 | no | |
ZTF24aadkwni | SN 2024aul | -18.0 | 59.8 ± 1.6 | 68.4 ± 1.6 | 0.84 ± 0.04 | -1.12 | no | |
ZTF21abjcjmc | SN 2021skn | -18.0 | 7.1 ± 0.0 | 55.7 ± 0.5 | 63.7 ± 0.5 | 2.05 ± 0.04 | -1.41 | yes |
ZTF23abascqa | SN 2023rbk | -18.0 | 71.9 ± 1.3 | 85.9 ± 1.3 | 1.28 ± 0.04 | -2.02 | no | |
ZTF24aaemydm | SN 2024chx | -18.0 | 64.3 ± 5.6 | 70.9 ± 5.6 | 0.92 ± 0.09 | -1.40 | no | |
ZTF24aajxppf | SN 2024grw | -18.0 | 52.7 ± 4.6 | 64.4 ± 4.6 | 1.37 ± 0.13 | -1.32 | no | |
ZTF21aaqugxm | SN 2021hdt | -18.1 | 49.1 ± 8.9 | 60.9 ± 8.9 | 1.34 ± 0.25 | -1.07 | no | |
ZTF21ablvzhp | SN 2021tiq | -18.1 | 85.9 ± 4.1 | 98.1 ± 4.1 | 0.56 ± 0.04 | -1.37 | yes | |
ZTF21abgilzj | AT 2021qcr | -18.1 | 16.2 ± 1.0 | 66.8 ± 1.1 | 93.0 ± 1.1 | 1.51 ± 0.04 | -0.92 | yes |
ZTF21aapkcmr | AT 2021fnj | -18.1 | 54.1 ± 4.6 | 59.1 ± 4.6 | 1.74 ± 0.15 | -1.34 | yes | |
ZTF19abqrhvt | SN 2019nyk | -18.1 | 47.4 ± 2.0 | 60.1 ± 2.0 | 1.54 ± 0.08 | -1.47 | yes | |
ZTF21abnudtb | SN 2021txr | -18.1 | 71.6 ± 1.0 | 80.5 ± 1.0 | 0.63 ± 0.03 | -1.16 | no | |
ZTF23aaphnyz | SN 2023lkw | -18.1 | 77.5 ± 10.7 | 87.7 ± 10.7 | 0.46 ± 0.07 | -0.95 | yes | |
ZTF22aaolwsd | SN 2022mxv | -18.1 | 12.6 ± 0.2 | 104.8 ± 20.0 | 122.6 ± 20.0 | 1.35 ± 0.26 | -0.87 | yes |
ZTF19aarykkb | SN 2019dzk | -18.2 | 73.9 ± 2.8 | 82.0 ± 2.8 | 0.3 ± 0.03 | -0.62 | yes | |
ZTF22aativsd | SN 2022ovb | -18.2 | 11.6 ± 0.2 | 71.2 ± 1.9 | 84.3 ± 1.9 | 0.84 ± 0.04 | -0.94 | no |
ZTF19abqrhvy | SN 2019odf | -18.2 | 14.0 ± 0.3 | 93.3 ± 8.4 | 107.8 ± 8.4 | 0.71 ± 0.07 | -1.14 | yes |
ZTF19abbwfgp | SN 2019ikb | -18.2 | 11.9 ± 0.3 | 79.1 ± 2.7 | 91.5 ± 2.7 | 1.05 ± 0.04 | -2.06 | yes |
ZTF22aavobvq | SN 2022prv | -18.3 | 10.6 ± 0.2 | 62.8 ± 12.0 | 86.5 ± 12.0 | 1.4 ± 0.27 | -2.56 | yes |
ZTF21aaipypa | SN 2021cgu | -18.3 | 13.2 ± 0.6 | 90.0 ± 13.1 | 106.5 ± 13.1 | 0.98 ± 0.14 | -0.45 | yes |
ZTF19acrcxri | SN 2019ult | -18.4 | 54.9 ± 1.6 | 65.3 ± 1.6 | 1.07 ± 0.05 | -1.33 | yes | |
ZTF23aaaatjn | SN 2023cf | -18.5 | 84.6 ± 5.7 | 90.1 ± 5.7 | 0.89 ± 0.06 | -1.92 | no | |
ZTF22aarycqo | SN 2022ojo | -19.3 | 80.8 ± 11.0 | 96.1 ± 11.0 | 1.42 ± 0.2 | -1.40 | no |
Parameter | Prior |
Initial radius [ (cm)] | |
Ejected mass [ ()] | |
Kinetic energy [ ( erg)] | |
Thermal energy [ ( erg)] | |
Thomson scattering opacity [ (cm2/g)] | |
Density power-law exponent | |
Ionization/recombination temperature [ (K)] | |
Nickel mass [ ()] | |
Gamma-leakage parameter [ (day2)] |
Parameter | Prior |
Explosion Energy [ erg] | |
Mass-Loss Rate [] | |
Wind Velocity [km s-1] | Fixed at 10 |
RCSM [ cm] | |
Wind Structure Parameter () |
ZTF name | IAU name | ||||
(M⊙) | (R⊙) | erg) | |||
ZTF21abrluay | SN 2021vfh | ||||
ZTF22abfxkdm | SN 2022ubb | ||||
ZTF23aasrcyv | SN 2023nlu | ||||
ZTF22abkhrkd | SN 2022wol | ||||
ZTF21aaagypx | SN 2021V | ||||
ZTF19aczkjxn | SN 2020ej | ||||
ZTF18aawpwlf | SN 2020hvn | ||||
ZTF22aafsqud | SN 2022hql | ||||
ZTF21acafqtj | SN 2021yok | ||||
ZTF21acpqqgu | SN 2021aewn | ||||
ZTF19aaqdkrm | SN 2019dod | ||||
ZTF19abqrhvt | SN 2019nyk | ||||
ZTF21aaipypa | SN 2021cgu | ||||
ZTF20acuhren | SN 2020abcq | ||||
ZTF24aaucrua | SN 2024nez | ||||
ZTF24abtczty | SN 2024abfl | ||||
ZTF22aapargp | SN 2022niw | ||||
ZTF23aaqknaw | SN 2023lzn | ||||
ZTF22abhsxph | SN 2022vyc | ||||
ZTF22abzqwmp | SN 2022adth | ||||
ZTF24aabsmvc | SN 2024ws | ||||
ZTF21aantsla | SN 2021ech | ||||
ZTF19aclobbu | SN 2019twk | ||||
ZTF21abnlhxs | SN 2021tyw | ||||
ZTF23aanymcl | SN 2023kzz | ||||
ZTF21abouuat | SN 2021ucg | ||||
ZTF19aakiyfn | AT 2019czo | ||||
ZTF19acrcxri | SN 2019ult | ||||
ZTF21aakvroo | SN 2021cwe | ||||
ZTF19acifduk | AT 2019toc | ||||
ZTF22absqhkw | SN 2022zkc | ||||
ZTF18abzrgim | SN 2018gvt | ||||
ZTF21aapliyn | SN 2021foj | ||||
ZTF21abjcjmc | SN 2021skn | ||||
ZTF22abyohff | SN 2022acrl | ||||
ZTF19aazudta | SN 2019hqm | ||||
ZTF19aamwhat | SN 2019bzd | ||||
ZTF19aazfvhh | SN 2019hpv | ||||
ZTF22abyokkf | SN 2022acri | ||||
ZTF24aaabbse | SN 2023achj | ||||
ZTF19abqrhvy | SN 2019odf | ||||
ZTF18aaszvfn | SN 2021iaw | ||||
ZTF23aaxadel | SN 2023pbg | ||||
ZTF22abfwxtr | SN 2022udq | ||||
ZTF21aaobkmg | SN 2021eui | ||||
ZTF24aabpzuz | SN 2024vs | ||||
ZTF19aawgxdn | SN 2019gmh | ||||
ZTF19aarykkb | SN 2019dzk |
ZTF | IAU | ||||
(M⊙) | (R⊙) | erg) | |||
ZTF22aazmrpx | SN 2022raj | ||||
ZTF21abnudtb | SN 2021txr | ||||
ZTF24aafqzur | SN 2024daa | ||||
ZTF19aamggnw | SN 2019bkp | ||||
ZTF22aavobvq | SN 2022prv | ||||
ZTF21ablvzhp | SN 2021tiq | ||||
ZTF23aaaatjn | SN 2023cf | ||||
ZTF21acgunkr | SN 2021aaxs | ||||
ZTF21abgilzj | AT 2021qcr | ||||
ZTF24aabppgn | SN 2024wp | ||||
ZTF20acmaaan | SN 2020xyk | ||||
ZTF23aackjhs | SN 1995al | ||||
ZTF22abtjefa | SN 2022aaad | ||||
ZTF24aaplfjd | SN 2024jxm | ||||
ZTF19aamhmsx | SN 2019bph | ||||
ZTF23aaphnyz | SN 2023lkw | ||||
ZTF23abnogui | SN 2023wcr | ||||
ZTF24aajxppf | SN 2024grw | ||||
ZTF24aaejecr | SN 2024btj | ||||
ZTF19aailepg | SN 2019amt | ||||
ZTF22aaolwsd | SN 2022mxv | ||||
ZTF19abqrhvt | SN 2019nyk | ||||
ZTF21aapkcmr | AT 2021fnj | ||||
ZTF21aaqugxm | SN 2021hdt | ||||
ZTF23abhzfww | SN 2023twg | ||||
ZTF24aaucrua | SN 2024nez | ||||
ZTF22abnujbv | SN 2022xus | ||||
ZTF22aativsd | SN 2022ovb | ||||
ZTF19acewuwn | SN 2019ssl | ||||
ZTF23aailjjs | SN 2023hcp | ||||
ZTF20aapchqy | SN 2020cxd | ||||
ZTF24aaemydm | SN 2024chx | ||||
ZTF21acgrrnl | SN 2021aayf | ||||
ZTF23abaxtlq | SN 2023rix | ||||
ZTF21aaeqwov | AT 2021htp | ||||
ZTF24aagupsf | SN 2024egd | ||||
ZTF22aavbfhz | AT 2022phi | ||||
ZTF19aauqwna | SN 2019fem | ||||
ZTF19acytcsg | SN 2019wvz | ||||
ZTF19aazyvub | SN 2019hnl | ||||
ZTF19aaniore | SN 2019ceg | ||||
ZTF22abkbjsb | SN 2022vym | ||||
ZTF19actnyae | SN 2019vdm | ||||
ZTF18abjndhh | SN 2018ecj | ||||
ZTF24aadkwni | SN 2024aul | ||||
ZTF19abajxet | SN 2019hyk | ||||
ZTF21abfiuqf | SN 2021pla | ||||
ZTF24abmkros | SN 2024xkd |
ZTF | IAU | ||||
(M⊙) | (R⊙) | erg) | |||
ZTF22aakdqqg | SN 2022kad | ||||
ZTF21aafepon | SN 2021ass | ||||
ZTF22abyivoq | SN 2022acko | ||||
ZTF21abvcxel | SN 2021wvw | ||||
ZTF23abbtkrv | SN 2023rvo | ||||
ZTF19aakzash | SN 2019bir | ||||
ZTF19abbnamr | SN 2019iex | ||||
ZTF20abeohfn | SN 2020mjm | ||||
ZTF19acbwejj | SN 2019upq | ||||
ZTF19abbwfgp | SN 2019ikb | ||||
ZTF22aarycqo | SN 2022ojo | ||||
ZTF21abhhrpj | SN 2021qiu | ||||
ZTF22abssiet | SN 2022zmb | ||||
ZTF22abtspsw | SN 2022aagp | ||||
ZTF21aavhnpk | SN 2021jsf | ||||
ZTF23abascqa | SN 2023rbk | ||||
ZTF19abwztsb | SN 2019pjs | ||||
ZTF22aaywnyg | SN 2022pru | ||||
ZTF19aanlekq | SN 2019cct | ||||
ZTF19actnwtn | SN 2019vdl | ||||
ZTF21aabygea | SN 2021os | ||||
ZTF23abmoxlu | SN 2023vog | ||||
ZTF22aakdbia | SN 2022jzc | ||||
ZTF19aadnxnl | SN 2019va | ||||
ZTF18adbacau | SN 2019aas | ||||
ZTF21aafkwtk | SN 2021apg | ||||
ZTF19aanrrqu | SN 2019clp | ||||
ZTF23aasbvab | SN 2023ngy | ||||
ZTF18aatyqds | SN 2018btl | ||||
ZTF21aardvtn | AT 2021htp | ||||
ZTF19acftfav | SN 2019ssi | ||||
ZTF19aapafit | SN 2019cvz | ||||
ZTF19acbvisk | SN 2019rms | ||||
ZTF24aarvbxj | SN 2024lby | ||||
ZTF21aagtqna | SN 2021brb | ||||
ZTF22aasojye | SN 2022omr | ||||
ZTF21aanzcuj | SN 2021enz | ||||
ZTF22abfavpu | SN 2022tmb | ||||
ZTF18aaxkqgy | SN 2018ccb | ||||
ZTF19aanhhal | SN 2019cec |
ZTF name | IAU name | ZAMS mass | Energy | log | |||
() | ) | ( cm) | ( erg) | ( yr-1) | |||
ZTF22abkhrkd | SN 2022wol | 10.24 0.43 | |||||
ZTF22abtjefa | SN 2022aaad | 11.03 0.63 | |||||
ZTF19abqrhvt | SN 2019nyk | 11.69 0.86 | |||||
ZTF19acewuwn | SN 2019ssl | 12.56 1.09 | |||||
ZTF21aaqugxm | SN 2021hdt | 14.82 1.01 | |||||
ZTF24aaasazz | SN 2024ov | 10.47 0.50 | |||||
ZTF20abyylgi | SN 2020svn | 13.38 1.42 | |||||
ZTF24aagupsf | SN 2024egd | 12.03 0.51 | |||||
ZTF21aanzcuj | SN 2021enz | 13.08 0.33 | |||||
ZTF20acveadu | SN 2020absk | 10.31 0.33 | |||||
ZTF19aamwhat | SN 2019bzd | 10.51 0.66 | |||||
ZTF19abbwfgp | SN 2019ikb | 12.37 1.15 | |||||
ZTF23aanymcl | SN 2023kzz | 11.63 0.73 | |||||
ZTF18aawpwlf | SN 2020hvn | 10.64 0.68 | |||||
ZTF18aatyqds | SN 2018btl | 12.31 0.45 | |||||
ZTF18aaxkqgy | SN 2018ccb | 12.06 1.13 | |||||
ZTF21acafqtj | SN 2021yok | 12.63 0.93 | |||||
ZTF20aaiovoy | SN 2020bad | 10.45 3.45 | |||||
ZTF21aafepon | SN 2021ass | 11.00 0.57 | |||||
ZTF23abhzfww | SN 2023twg | 11.96 0.64 | |||||
ZTF20abyzomt | SN 2020szs | 12.08 0.62 | |||||
ZTF24aaucrua | SN 2024nez | 9.86 0.49 | |||||
ZTF21aaobkmg | SN 2021eui | 10.54 0.48 | |||||
ZTF21aagtqna | SN 2021brb | 13.02 0.81 | |||||
ZTF19aadnxnl | SN 2019va | 14.05 0.25 | |||||
ZTF19aanrrqu | SN 2019clp | 12.11 0.94 | |||||
ZTF20aciojzg | AT 2020wip | 10.43 0.50 | |||||
ZTF22abssiet | SN 2022zmb | 10.53 0.88 | |||||
ZTF18aaqfqaa | SN 2018bek | 10.00 0.55 | |||||
ZTF19aailepg | SN 2019amt | 12.55 0.48 | |||||
ZTF23abaxtlq | SN 2023rix | 10.70 0.34 | |||||
ZTF21aapkcmr | AT 2021fnj | 11.94 1.00 | |||||
ZTF18aawgrxz | SN 2021lmp | 10.81 0.89 | |||||
ZTF21acgrrnl | SN 2021aayf | 13.89 2.66 | |||||
ZTF22abkbjsb | SN 2022vym | 10.95 0.65 | |||||
ZTF19aclobbu | SN 2019twk | 11.55 0.97 | |||||
ZTF22aafsqud | SN 2022hql | 11.93 0.89 | |||||
ZTF22abfavpu | SN 2022tmb | 14.35 2.06 | |||||
ZTF21abfiuqf | SN 2021pla | 10.50 1.47 | |||||
ZTF24aarvbxj | SN 2024lby | 11.73 1.00 | |||||
ZTF24aaemydm | SN 2024chx | 10.17 0.59 | |||||
ZTF22abtspsw | SN 2022aagp | 11.40 1.62 | |||||
ZTF20acfvprb | AT 2020thp | 12.12 0.49 | |||||
ZTF23aasbvab | SN 2023ngy | 11.80 0.76 | |||||
ZTF22aaevwec | SN 2022gwg | 11.03 1.23 | |||||
ZTF21acfajbc | SN 2021aalq | 10.77 0.66 | |||||
ZTF24aadkwni | SN 2024aul | 10.32 1.02 | |||||
ZTF21abouuat | SN 2021ucg | 16.65 1.01 |
ZTF name | IAU name | ZAMS mass | Energy | ||||
() | ) | ( cm) | ( erg) | (log yr-1) | |||
ZTF18aahfqbc | SN 2018aoq | 12.05 0.78 | |||||
ZTF21abhhrpj | SN 2021qiu | 15.71 2.99 | |||||
ZTF20actnuls | SN 2020aawo | 12.06 0.57 | |||||
ZTF22abnujbv | SN 2022xus | 10.57 1.34 | |||||
ZTF23abbtkrv | SN 2023rvo | 15.79 0.58 | |||||
ZTF19aaniore | SN 2019ceg | 12.19 1.24 | |||||
ZTF22aasojye | SN 2022omr | 13.16 0.51 | |||||
ZTF20aafckit | SN 2020uu | 10.24 0.99 | |||||
ZTF21aaagypx | SN 2021V | 14.23 1.12 | |||||
ZTF19acbwejj | SN 2019upq | 14.86 2.75 | |||||
ZTF23aackjhs | SN 1995al | 9.46 0.70 | |||||
ZTF21aaeqwov | AT 2021htp | 10.40 1.14 | |||||
ZTF19aanhhal | SN 2019cec | 14.22 3.80 | |||||
ZTF22aanrqje | SN 2022mji | 9.41 0.67 | |||||
ZTF21aafkwtk | SN 2021apg | 12.42 0.36 | |||||
ZTF24abmkros | SN 2024xkd | 10.82 1.04 | |||||
ZTF22aaolwsd | SN 2022mxv | 11.75 0.81 | |||||
ZTF24abtczty | SN 2024abfl | 10.71 0.45 | |||||
ZTF19abajxet | SN 2019hyk | 11.93 0.55 | |||||
ZTF24aaplfjd | SN 2024jxm | 10.39 0.34 | |||||
ZTF21aantsla | SN 2021ech | 11.06 0.57 | |||||
ZTF24aajxppf | SN 2024grw | 10.96 0.93 | |||||
ZTF19acftfav | SN 2019ssi | 10.47 0.37 | |||||
ZTF20actkmks | SN 2020aaun | 9.98 2.83 | |||||
ZTF21abnlhxs | SN 2021tyw | 14.43 1.29 | |||||
ZTF23aaphnyz | SN 2023lkw | 12.80 0.32 | |||||
ZTF19actnwtn | SN 2019vdl | 12.05 0.48 | |||||
ZTF18aadsuxd | SN 2020yui | 12.03 1.36 | |||||
ZTF22aativsd | SN 2022ovb | 12.08 0.40 | |||||
ZTF23aaqknaw | SN 2023lzn | 13.69 0.73 | |||||
ZTF24aabsmvc | SN 2024ws | 11.32 0.87 | |||||
ZTF22abvaetz | SN 2022aang | 10.21 0.66 | |||||
ZTF22ablnrcv | SN 2022xav | 9.28 0.68 | |||||
ZTF21aardvtn | AT 2021htp | 10.22 0.23 | |||||
ZTF22aauurbv | SN 2022pgf | 10.37 0.36 | |||||
ZTF19aazyvub | SN 2019hnl | 11.61 0.51 | |||||
ZTF24aaejecr | SN 2024btj | 10.54 0.83 | |||||
ZTF22aakdqqg | SN 2022kad | 11.95 0.76 | |||||
ZTF22aarycqo | SN 2022ojo | 13.18 1.81 | |||||
ZTF23abascqa | SN 2023rbk | 10.28 1.02 | |||||
ZTF23aasrcyv | SN 2023nlu | 11.00 1.15 | |||||
ZTF23aailjjs | SN 2023hcp | 11.81 0.46 | |||||
ZTF21aabygea | SN 2021os | 12.50 1.34 | |||||
ZTF20acvjagm | SN 2020abtf | 12.73 0.36 | |||||
ZTF19actnyae | SN 2019vdm | 10.50 0.99 | |||||
ZTF21ablvzhp | SN 2021tiq | 10.09 0.68 | |||||
ZTF21aavhnpk | SN 2021jsf | 9.59 1.34 | |||||
ZTF22aaudjgc | SN 2022pfx | 10.45 0.71 |
ZTF name | IAU name | ZAMS mass | Energy | ||||
() | ) | ( cm) | ( erg) | (log yr-1) | |||
ZTF22abfxkdm | SN 2022ubb | 10.62 0.43 | |||||
ZTF22aavbfhz | AT 2022phi | 10.64 1.08 | |||||
ZTF22aakdbia | SN 2022jzc | 9.92 0.15 | |||||
ZTF22aazmrpx | SN 2022raj | 10.94 0.50 | |||||
ZTF20acfkzcg | AT 2021htp | 11.40 0.77 | |||||
ZTF21abrluay | SN 2021vfh | 12.92 1.37 | |||||
ZTF19acytcsg | SN 2019wvz | 15.04 0.64 | |||||
ZTF19abwztsb | SN 2019pjs | 10.72 0.76 | |||||
ZTF22aaywnyg | SN 2022pru | 10.86 0.89 | |||||
ZTF19abbnamr | SN 2019iex | 14.16 0.73 | |||||
ZTF23abbsxzs | SN 2023rtq | 9.84 1.38 | |||||
ZTF22abzqwmp | SN 2022adth | 11.95 0.56 | |||||
ZTF20acitoie | SN 2020wgp | 10.74 0.97 | |||||
ZTF19aapafit | SN 2019cvz | 14.95 1.07 | |||||
ZTF20acbjgoa | SN 2020acac | 10.48 0.64 | |||||
ZTF21abvcxel | SN 2021wvw | 9.56 0.57 | |||||
ZTF20aapchqy | SN 2020cxd | 9.90 0.08 | |||||
ZTF18aaszvfn | SN 2021iaw | 12.27 0.62 | |||||
ZTF20acuhren | SN 2020abcq | 10.28 0.97 | |||||
ZTF23abnogui | SN 2023wcr | 10.75 1.02 | |||||
ZTF22abyivoq | SN 2022acko | 10.97 0.25 | |||||
ZTF22abhsxph | SN 2022vyc | 12.20 0.91 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2019env | 2019-06-25 | 51 | P60 | |
SN 2018bjo | 2018-06-12 | 56 | P200 | |
SN 2018cbb | 2018-05-17 | 45 | P200 | |
SN 2018bdv | 2018-05-11 | 16 | P60 | |
SN 2024iwm | 2024-06-01 | 18 | SEDM | |
SN 2024iwm | 2024-06-06 | 23 | Deveny+LMI | |
SN 2024iwm | 2024-06-07 | 24 | KAST | |
SN 2021iaw | 2021-05-06 | 33 | KAST | |
SN 2018btl | 2018-07-16 | 63 | P60 | |
SN 2021lmp | 2021-05-29 | 24 | SEDM | |
SN 2021lmp | 2021-05-31 | 25 | DBSP | |
SN 2020hvn | 2020-05-07 | 17 | P60 | |
SN 2018lrq | 2018-06-21 | 69 | P200 | |
SN 2018cqp | 2018-07-02 | 18 | P60 | |
SN 2018cwa | 2018-08-04 | 43 | P200 | |
SN 2018dzc | 2018-07-30 | 20 | P60 | |
SN 2018dzc | 2018-07-30 | 20 | P60 | |
SN 2018ecj | 2018-12-01 | 129 | Keck1 | |
SN 2018dht | 2018-08-04 | 24 | P60 | |
SN 2018dht | 2018-08-04 | 24 | P200 | |
SN 2018dht | 2018-08-04 | 24 | P60 | |
AT 2018lrz | 2018-09-12 | 25 | P200 | |
SN 2018fif | 2018-09-04 | 14 | NOT | |
SN 2018fif | 2018-09-25 | 35 | P60 | |
SN 2018fif | 2018-09-25 | 35 | P60 | |
SN 2018fif | 2018-11-03 | 73 | P60 | |
SN 2018fif | 2018-11-14 | 84 | P60 | |
SN 2018fif | 2018-11-19 | 89 | P60 | |
SN 2018fif | 2018-11-26 | 96 | P60 | |
SN 2018fif | 2018-12-04 | 103 | P60 | |
AT 2018lsa | 2018-09-12 | 17 | P200 | |
SN 2024kan | 2024-06-07 | 13 | SEDM | |
SN 2024kan | 2024-06-15 | 21 | DBSP | |
SN 2018gvt | 2018-12-05 | 69 | P60 | |
SN 2018gvt | 2018-12-23 | 86 | P60 | |
SN 2018hle | 2018-12-24 | 57 | P60 | |
AT 2018joy | 2018-12-01 | 24 | Keck1 | |
AT 2018jwq | 2018-12-01 | 14 | Keck1 | |
AT 2018kds | 2019-01-26 | 58 | P200 | |
SN 2018hwm | 2018-12-14 | 33 | P200 | |
SN 2018hwm | 2018-12-17 | 36 | P60 | |
SN 2018kkv | 2019-01-26 | 37 | P200 | |
SN 2018lev | 2019-01-03 | 13 | P60 | |
SN 2019va | 2019-03-13 | 56 | P60 | |
SN 2019vb | 2019-01-26 | 10 | P200 | |
SN 2019amt | 2019-03-15 | 37 | P200 | |
AT 2019aor | 2019-03-19 | 83 | P60 | |
AT 2019aor | 2019-04-06 | 100 | P60 | |
AT 2019txj | 2019-03-15 | 13 | P200 | |
AT 2019txj | 2019-04-03 | 32 | Keck1 | |
SN 2019dvw | 2019-04-13 | 51 | P200 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2019dvw | 2019-04-24 | 62 | P200 | |
SN 2019dvd | 2019-05-02 | 12 | P60 | |
AT 2019esn | 2019-05-13 | 13 | P200 | |
SN 2019gss | 2019-06-26 | 29 | P200 | |
SN 2019gmh | 2019-07-03 | 32 | P60 | |
SN 2019gmh | 2019-08-15 | 73 | P60 | |
SN 2019gmh | 2019-09-09 | 98 | P60 | |
SN 2019gmh | 2019-09-18 | 106 | P60 | |
SN 2019gmh | 2019-09-18 | 106 | P60 | |
SN 2019gmh | 2019-09-19 | 107 | P60 | |
SN 2019gmh | 2019-09-19 | 107 | P60 | |
SN 2019gmh | 2019-09-24 | 112 | P60 | |
SN 2019gmh | 2019-09-30 | 118 | P60 | |
SN 2019hrb | 2019-07-08 | 29 | P60 | |
SN 2019hnl | 2019-07-03 | 17 | P60 | |
AT 2019khq | 2019-06-26 | 14 | P200 | |
SN 2019iex | 2019-07-07 | 12 | P60 | |
SN 2019iex | 2019-10-21 | 116 | P60 | |
SN 2019ikb | 2019-07-14 | 17 | P60 | |
SN 2019ikb | 2019-08-04 | 37 | P60 | |
SN 2019ikb | 2019-10-08 | 101 | P60 | |
SN 2019ikb | 2019-10-15 | 107 | P60 | |
SN 2019ikb | 2019-10-15 | 107 | P60 | |
AT 2019tti | 2019-08-31 | 80 | P200 | |
SN 2019mkr | 2019-08-10 | 17 | P60 | |
SN 2019mor | 2019-08-19 | 20 | P60 | |
SN 2019mor | 2019-08-28 | 29 | P60 | |
SN 2019mor | 2019-08-28 | 29 | P60 | |
SN 2019mor | 2019-09-02 | 34 | P200 | |
AT 2019tya | 2019-08-10 | 11 | P200 | |
AT 2019ttl | 2019-08-04 | 16 | P200 | |
SN 2019nvm | 2019-09-01 | 13 | UH88 | |
SN 2019nvm | 2019-09-06 | 18 | NOT | |
SN 2019nvm | 2019-11-02 | 74 | P60 | |
SN 2019nvm | 2019-11-13 | 85 | P60 | |
SN 2019nvm | 2019-11-14 | 86 | P60 | |
SN 2019nvm | 2020-01-29 | 160 | NOT | |
SN 2019nyk | 2019-09-19 | 31 | P60 | |
SN 2019odf | 2019-09-22 | 32 | P60 | |
SN 2019qba | 2019-09-21 | 11 | P60 | |
SN 2019qba | 2019-09-21 | 11 | P60 | |
AT 2019qiq | 2019-09-28 | 18 | Keck1 | |
SN 2019rho | 2019-10-11 | 13 | P60 | |
SN 2019upq | 2020-01-05 | 97 | P60 | |
SN 2019ssi | 2019-11-18 | 31 | P60 | |
SN 2019twk | 2019-11-13 | 13 | P60 | |
SN 2019twk | 2019-11-24 | 23 | P60 | |
AT 2019ubr | 2019-11-05 | 22 | P200 | |
SN 2019ult | 2019-11-24 | 16 | P200 | |
SN 2019ust | 2019-11-23 | 12 | P60 | |
SN 2019ust | 2019-11-24 | 13 | P200 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2019ust | 2019-11-24 | 13 | P200 | |
SN 2019ust | 2019-11-26 | 14 | P60 | |
SN 2019ust | 2019-12-10 | 28 | P60 | |
SN 2019ust | 2020-01-02 | 51 | P60 | |
SN 2019vdl | 2019-12-16 | 29 | P60 | |
SN 2019vdl | 2019-12-18 | 31 | P60 | |
SN 2019vdm | 2019-12-19 | 31 | P60 | |
SN 2019vdm | 2019-12-23 | 35 | Keck1 | |
SN 2019vjl | 2019-12-11 | 19 | P60 | |
SN 2019vsr | 2019-12-18 | 18 | P60 | |
SN 2019vsr | 2019-12-19 | 19 | P60 | |
SN 2019vsr | 2019-12-23 | 23 | LT | |
SN 2019vsr | 2019-12-23 | 23 | LT | |
SN 2019wbd | 2019-12-13 | 14 | P60 | |
SN 2019wbd | 2019-12-21 | 22 | P200 | |
SN 2019wbd | 2019-12-27 | 28 | LT | |
SN 2019xbm | 2020-01-07 | 31 | P60 | |
SN 2019xbm | 2020-01-27 | 51 | P200 | |
SN 2019wvz | 2019-12-28 | 12 | LT | |
SN 2019xgi | 2019-12-31 | 14 | P60 | |
SN 2019xgi | 2019-12-31 | 14 | P60 | |
SN 2019xgi | 2020-01-01 | 15 | P200 | |
SN 2019xgi | 2020-01-01 | 15 | P200 | |
SN 2019xvo | 2020-01-12 | 18 | P60 | |
SN 2019xvo | 2020-01-12 | 18 | P60 | |
SN 2019xvo | 2020-01-27 | 33 | P200 | |
AT 2020og | 2020-01-27 | 26 | P200 | |
AT 2020og | 2020-02-14 | 43 | P200 | |
SN 2020uu | 2020-01-27 | 18 | P200 | |
SN 2020aqb | 2020-02-13 | 24 | P60 | |
SN 2020aqy | 2020-02-02 | 12 | P60 | |
SN 2020alg | 2020-01-31 | 11 | P60 | |
SN 2020alg | 2020-02-08 | 19 | LT | |
SN 2020awg | 2020-02-07 | 21 | P60 | |
AT 2020bow | 2020-02-14 | 79 | P200 | |
SN 2020cxd | 2020-05-22 | 92 | P60 | |
SN 2020cxd | 2020-06-25 | 126 | LT | |
SN 2020cxd | 2020-06-26 | 127 | P60 | |
SN 2020cxd | 2020-07-01 | 132 | NOT | |
SN 2020cxd | 2020-07-28 | 159 | Gemini | |
SN 2020cxd | 2020-10-19 | 241 | Keck1 | |
SN 2020edr | 2020-03-06 | 16 | P60 | |
SN 2020edr | 2020-06-12 | 112 | P60 | |
SN 2020edr | 2020-06-25 | 125 | P200 | |
SN 2020edr | 2020-06-25 | 125 | P200 | |
SN 2020djn | 2020-03-04 | 10 | P60 | |
SN 2020djn | 2020-03-04 | 10 | P60 | |
SN 2020drl | 2020-03-31 | 34 | P60 | |
SN 2020drl | 2020-04-15 | 49 | P60 | |
SN 2020drl | 2020-04-15 | 49 | P60 | |
AT 2020dzo | 2020-03-22 | 21 | Keck1 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2020eaf | 2021-01-12 | 311 | LRIS | |
SN 2020fqv | 2020-04-15 | 16 | P60 | |
SN 2020gpe | 2020-05-29 | 48 | P60 | |
SN 2020ghv | 2020-04-23 | 12 | P60 | |
SN 2020hyb | 2020-05-18 | 27 | LT | |
SN 2020jfo | 2020-05-15 | 11 | P60 | |
SN 2020jfo | 2020-05-25 | 21 | P60 | |
SN 2020jfo | 2020-05-25 | 21 | KAST | |
SN 2020jfo | 2020-05-29 | 25 | KAST | |
SN 2020jfo | 2020-06-12 | 39 | P60 | |
SN 2020jfo | 2020-06-17 | 44 | P60 | |
SN 2020jfo | 2020-12-04 | 213 | P60 | |
SN 2020jfo | 2020-12-07 | 216 | ALFOSC | |
SN 2020jfo | 2020-12-07 | 216 | NOT | |
SN 2020jfo | 2020-12-22 | 231 | SEDM | |
SN 2020jfo | 2021-01-16 | 255 | ALFOSC | |
SN 2020jfo | 2021-01-16 | 255 | NOT | |
SN 2020jfo | 2021-02-04 | 274 | SEDM | |
SN 2020jfo | 2021-02-09 | 279 | NOT | |
SN 2020jfo | 2021-02-09 | 279 | ALFOSC | |
SN 2020jfo | 2021-03-09 | 307 | ALFOSC | |
SN 2020jfo | 2021-03-09 | 307 | NOT | |
SN 2020jfo | 2021-04-20 | 349 | ALFOSC | |
SN 2020jfo | 2021-04-20 | 349 | NOT | |
SN 2020jjj | 2020-05-23 | 17 | P60 | |
SN 2020kay | 2020-05-26 | 11 | P60 | |
SN 2020lam | 2020-06-04 | 10 | P60 | |
SN 2020lam | 2020-06-06 | 12 | NOT | |
SN 2020lrr | 2020-06-20 | 17 | P60 | |
SN 2020lrr | 2020-06-23 | 20 | Keck1 | |
SN 2020nif | 2020-07-03 | 10 | P60 | |
SN 2020nif | 2020-07-04 | 11 | P60 | |
SN 2020nif | 2020-07-23 | 30 | LT | |
SN 2020nny | 2020-06-27 | 20 | P60 | |
SN 2020noe | 2020-07-16 | 26 | P200 | |
SN 2020noe | 2020-07-16 | 26 | P200 | |
SN 2020nyb | 2020-07-15 | 15 | P60 | |
SN 2020nyb | 2020-07-15 | 15 | P60 | |
SN 2020oco | 2020-07-17 | 18 | P60 | |
SN 2020oco | 2020-07-21 | 22 | P60 | |
SN 2020oco | 2020-07-21 | 22 | P60 | |
SN 2020oco | 2020-07-31 | 31 | NOT | |
SN 2020oco | 2020-07-31 | 31 | NOT | |
SN 2020oco | 2020-08-10 | 41 | P60 | |
SN 2020ofw | 2020-08-07 | 29 | NOT | |
SN 2020qmp | 2020-12-12 | 140 | DBSP | |
SN 2020qmp | 2021-01-08 | 167 | DBSP | |
SN 2020qmp | 2021-02-20 | 210 | DBSP | |
AT 2020rid | 2020-08-15 | 24 | NOT | |
AT 2020afdi | 2020-09-13 | 35 | NOT | |
SN 2020sje | 2020-09-20 | 22 | Keck1 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2020sje | 2020-10-11 | 42 | P60 | |
SN 2020svn | 2020-09-19 | 12 | P60 | |
SN 2020svn | 2020-09-20 | 13 | P60 | |
SN 2020svn | 2020-09-21 | 14 | P60 | |
SN 2020svn | 2020-09-22 | 15 | P60 | |
SN 2020szs | 2020-10-27 | 45 | P60 | |
SN 2020szs | 2020-10-27 | 45 | SEDM | |
SN 2020tdm | 2020-09-27 | 15 | P60 | |
AT 2020ugt | 2020-09-28 | 11 | P60 | |
AT 2020ugt | 2020-10-21 | 33 | P200 | |
AT 2020ugt | 2020-10-21 | 33 | DBSP | |
SN 2020twk | 2020-10-03 | 13 | LT | |
SN 2020twk | 2020-10-03 | 13 | P60 | |
SN 2020twk | 2020-10-04 | 14 | P60 | |
SN 2020uim | 2020-10-11 | 14 | P60 | |
SN 2020uim | 2020-10-11 | 14 | P60 | |
SN 2020uim | 2020-12-12 | 75 | SEDM | |
SN 2020ult | 2020-10-21 | 25 | DBSP | |
SN 2020ult | 2020-10-21 | 25 | P200 | |
AT 2021htp | 2020-11-11 | 44 | P60 | |
AT 2020thp | 2020-10-10 | 30 | P60 | |
AT 2020thp | 2020-10-21 | 40 | P200 | |
AT 2020thp | 2020-10-21 | 40 | DBSP | |
SN 2020uqx | 2020-10-12 | 12 | P60 | |
SN 2020uqx | 2020-10-21 | 20 | DBSP | |
SN 2020uqx | 2020-10-21 | 20 | P200 | |
SN 2020uqx | 2020-10-22 | 21 | P60 | |
SN 2020vef | 2020-10-15 | 11 | P60 | |
SN 2020vef | 2020-11-12 | 39 | P60 | |
AT 2020wip | 2020-10-20 | 19 | Keck1 | |
AT 2020wip | 2020-12-11 | 70 | SEDM | |
SN 2020wgp | 2020-10-23 | 12 | P60 | |
SN 2020wgs | 2020-10-23 | 12 | P60 | |
SN 2020wgs | 2020-10-28 | 17 | P60 | |
SN 2020wgs | 2020-10-30 | 19 | LT | |
SN 2020wgs | 2020-10-30 | 19 | LT | |
SN 2020xhs | 2020-12-16 | 58 | SEDM | |
SN 2020aasd | 2020-12-07 | 15 | LT | |
SN 2020aaun | 2020-12-13 | 21 | SEDM | |
SN 2020abah | 2020-12-12 | 23 | LRIS | |
SN 2020aawo | 2020-12-04 | 11 | SEDM | |
SN 2020abbl | 2020-12-12 | 20 | LRIS | |
SN 2020abbi | 2020-12-18 | 26 | SEDM | |
SN 2020abcq | 2020-12-10 | 12 | SNIFS | |
SN 2020abcq | 2020-12-12 | 14 | SEDM | |
SN 2020absk | 2020-12-18 | 14 | LT | |
SN 2020absk | 2021-02-07 | 64 | SEDM | |
SN 2020absj | 2020-12-12 | 79 | SEDM | |
SN 2020abtf | 2020-12-11 | 12 | EFOSC2 | |
SN 2020abtf | 2020-12-16 | 17 | SEDM | |
SN 2020abtf | 8591-12-03 | 2366876 | EFOSC2 |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2020acjg | 2021-11-06 | 140 | DBSP | |
SN 2020acjg | 2022-02-05 | 228 | LRIS | |
SN 2020acjg | 2022-03-04 | 255 | LRIS | |
SN 2020adhh | 2021-01-09 | 21 | DIS | |
SN 2020aeqx | 2021-01-12 | 15 | LRIS | |
SN 2021V | 2021-01-12 | 12 | SEDM | |
SN 2021V | 2021-02-08 | 39 | SEDM | |
SN 2021aek | 2021-01-26 | 14 | EFOSC2 | |
SN 2021aek | 2021-02-04 | 23 | EFOSC2 | |
AT 2021htp | 2021-02-20 | 143 | DBSP | |
SN 2021brb | 2021-02-04 | 10 | SEDM | |
SN 2021brb | 2021-02-19 | 25 | SEDM | |
SN 2021cgu | 2021-03-09 | 29 | SEDM | |
SN 2021cwe | 2021-03-07 | 22 | SEDM | |
SN 2021cwe | 2021-04-07 | 52 | SEDM | |
SN 2021dvk | 2021-04-09 | 42 | DBSP | |
SN 2021enz | 2021-03-17 | 14 | SEDM | |
SN 2021eui | 2021-03-17 | 16 | SPRAT | |
SN 2021eui | 2021-05-31 | 90 | SEDM | |
SN 2021gmj | 2021-04-05 | 17 | SEDM | |
SN 2021gvm | 2021-04-03 | 14 | SPRAT | |
SN 2021hac | 2021-04-08 | 18 | SEDM | |
SN 2021hac | 2021-05-13 | 52 | NOT | |
SN 2021hac | 2021-05-13 | 52 | ALFOSC | |
SN 2021mju | 2021-05-31 | 71 | DBSP | |
SN 2021jsf | 2021-04-20 | 11 | SEDM | |
SN 2021jsf | 2021-04-20 | 11 | SNIFS | |
SN 2021jsf | 2021-06-09 | 59 | SEDM | |
SN 2021jsf | 2021-06-12 | 62 | DBSP | |
AT 2020ghq | 2021-05-24 | 10 | SEDM | |
SN 2021nli | 2021-05-31 | 10 | SEDM | |
SN 2021nli | 2021-06-04 | 14 | DBSP | |
AT 2014gz | 2021-06-19 | 24 | DIS | |
AT 2021pkh | 2021-07-10 | 32 | DBSP | |
SN 2021pqj | 2021-06-12 | 10 | DBSP | |
SN 2021pnh | 2021-07-09 | 29 | LRIS | |
SN 2021qiu | 2021-06-28 | 12 | SPRAT | |
SN 2021qiu | 2021-07-05 | 19 | SEDM | |
SN 2021qzi | 2021-07-05 | 10 | SEDM | |
SN 2021skm | 2021-08-01 | 53 | DBSP | |
SN 2021svy | 2021-07-30 | 22 | SEDM | |
SN 2021svy | 2021-08-04 | 27 | DBSP | |
SN 2021tyw | 2021-08-03 | 11 | EFOSC2 | |
SN 2021tyw | 2021-08-03 | 11 | EFOSC2 | |
SN 2021tyw | 2021-08-05 | 13 | EFOSC2 | |
SN 2021tyw | 2021-08-05 | 13 | EFOSC2 | |
SN 2021tyw | 2021-08-06 | 14 | EFOSC2 | |
SN 2021tyw | 2021-08-06 | 14 | EFOSC2 | |
SN 2021tyw | 2021-08-12 | 20 | EFOSC2 | |
SN 2021tyw | 2021-08-12 | 20 | KAST | |
SN 2021tyw | 2021-08-12 | 20 | SEDM |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2021tyw | 2021-08-12 | 20 | EFOSC2 | |
SN 2021tyw | 2021-08-17 | 25 | SEDM | |
SN 2021tyw | 2021-08-21 | 29 | SEDM | |
SN 2021tyw | 2021-08-30 | 38 | SEDM | |
SN 2021tyw | 2021-09-13 | 51 | EFOSC2 | |
SN 2021tyw | 2021-09-13 | 51 | EFOSC2 | |
SN 2021tyw | 2021-11-09 | 108 | SEDM | |
SN 2021ucg | 2022-04-27 | 271 | LRIS | |
SN 2021vfh | 2021-08-21 | 14 | SEDM | |
SN 2021vfh | 2021-08-22 | 15 | SPRAT | |
SN 2021wun | 2021-09-06 | 36 | SEDM | |
SN 2021wun | 2022-02-02 | 182 | DBSP | |
SN 2021wvw | 2021-09-25 | 32 | NIRES | |
SN 2021xat | 2021-09-08 | 16 | SEDM | |
SN 2021ywg | 2021-09-25 | 12 | NIRES | |
SN 2021ywg | 2021-09-27 | 14 | DBSP | |
SN 2021yyg | 2021-09-27 | 13 | DBSP | |
SN 2021zco | 2021-09-27 | 11 | DBSP | |
SN 2021zco | 2021-09-29 | 13 | SEDM | |
SN 2021zex | 2021-12-19 | 89 | SEDM | |
SN 2021zzi | 2021-10-13 | 15 | DBSP | |
SN 2021aatd | 2021-10-16 | 10 | SEDM | |
SN 2021aatd | 2021-12-06 | 60 | SEDM | |
SN 2021aatd | 2021-12-12 | 66 | SEDM | |
SN 2021aatd | 2021-12-18 | 72 | SEDM | |
SN 2021aatd | 2022-01-12 | 97 | SEDM | |
SN 2021aatd | 2022-01-15 | 100 | SEDM | |
SN 2021aatd | 2022-01-25 | 110 | SEDM | |
SN 2021aatd | 2022-01-25 | 110 | SEDM | |
SN 2021aatd | 2022-01-28 | 112 | SEDM | |
SN 2021aatd | 2022-01-31 | 115 | LRIS | |
SN 2021aatd | 2022-02-08 | 123 | SEDM | |
SN 2021aatd | 2022-02-12 | 127 | SEDM | |
SN 2021aatd | 2022-08-23 | 316 | LRIS | |
SN 2021aatd | 2022-08-23 | 316 | LRIS | |
SN 2021aatd | 2022-10-23 | 376 | LRIS | |
SN 2021aatd | 2022-10-23 | 376 | LRIS | |
SN 2021aaxs | 2021-10-15 | 11 | SEDM | |
SN 2021aaxs | 2021-10-21 | 17 | SEDM | |
SN 2021abpd | 2021-10-25 | 12 | SEDM | |
SN 2021abqs | 2021-10-27 | 11 | SEDM | |
SN 2021achr | 2021-10-31 | 20 | SEDM | |
SN 2021achr | 2021-11-06 | 26 | DBSP | |
SN 2021adxd | 2021-11-20 | 15 | SEDM | |
SN 2021adxd | 2021-11-29 | 24 | SEDM | |
SN 2021adxd | 2021-12-01 | 26 | DBSP | |
SN 2021aewn | 2021-12-01 | 14 | DBSP | |
SN 2021aewn | 2021-12-03 | 16 | SEDM | |
SN 2021aewn | 2021-12-12 | 25 | SEDM | |
SN 2021afud | 2021-12-08 | 10 | SEDM | |
SN 2021afud | 2021-12-11 | 13 | SEDM |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2021afud | 2022-01-12 | 44 | DBSP | |
SN 2022cru | 2022-04-27 | 118 | LRIS | |
SN 2022ces | 2022-03-02 | 25 | SEDM | |
SN 2022gwg | 2022-04-19 | 12 | SEDM | |
SN 2022hql | 2022-04-26 | 12 | SEDM | |
SN 2022hss | 2022-05-02 | 16 | SEDM | |
SN 2022iob | 2022-05-23 | 37 | DBSP | |
SN 2022iyl | 2022-06-08 | 42 | DBSP | |
SN 2022joe | 2022-05-23 | 17 | DBSP | |
SN 2022jux | 2022-05-23 | 11 | SEDM | |
SN 2022jzc | 2022-05-31 | 16 | SEDM | |
SN 2022jzc | 2022-06-07 | 23 | SEDM | |
SN 2022jzc | 2022-06-08 | 24 | DBSP | |
SN 2022jzc | 2022-06-21 | 37 | SEDM | |
SN 2022jzc | 2022-06-27 | 43 | SEDM | |
SN 2022jzc | 2022-07-01 | 47 | SEDM | |
SN 2022jzc | 2022-07-04 | 50 | SEDM | |
SN 2022jzc | 2022-07-07 | 53 | ALFOSC | |
SN 2022jzc | 2022-07-15 | 61 | SEDM | |
SN 2022jzc | 2022-07-20 | 66 | ALFOSC | |
SN 2022mxv | 2022-08-21 | 64 | SEDM | |
AT 2022ohx | 2022-07-27 | 25 | LRIS | |
SN 2022oor | 2022-07-29 | 21 | DBSP | |
SN 2022omr | 2022-07-21 | 14 | DBSP | |
SN 2022omr | 2022-08-03 | 26 | LRIS | |
SN 2022ovb | 2022-07-24 | 10 | SEDM | |
SN 2022paf | 2022-07-29 | 14 | DBSP | |
SN 2022oyp | 2022-07-29 | 13 | DBSP | |
SN 2022oyp | 2022-08-23 | 37 | SEDM | |
AT 2022phi | 2022-08-26 | 37 | DBSP | |
SN 2022prv | 2022-08-11 | 20 | SEDM | |
SN 2022prv | 2022-08-12 | 21 | SPRAT | |
SN 2022prv | 2022-08-19 | 28 | SPRAT | |
SN 2022prv | 2022-08-24 | 33 | SEDM | |
SN 2022prv | 2022-08-25 | 34 | SEDM | |
SN 2022prv | 2022-09-01 | 41 | SEDM | |
SN 2022prv | 2022-09-07 | 47 | SEDM | |
SN 2022raj | 2022-08-20 | 12 | DBSP | |
SN 2022raj | 2023-10-07 | 420 | LRIS | |
SN 2022rqg | 2022-08-23 | 12 | LRIS | |
SN 2022udq | 2022-10-02 | 19 | DBSP | |
SN 2022udq | 2022-10-06 | 23 | SEDM | |
SN 2022vyc | 2022-10-04 | 11 | SEDM | |
SN 2022wbr | 2022-10-02 | 16 | DBSP | |
SN 2022wbr | 2022-11-17 | 61 | LRIS | |
SN 2022wbr | 2023-01-17 | 121 | LRIS | |
SN 2022wbr | 2023-11-20 | 422 | LRIS | |
SN 2022wol | 2022-11-26 | 54 | DBSP | |
SN 2022wol | 2022-12-16 | 74 | DBSP | |
SN 2022xav | 2022-10-21 | 16 | SEDM | |
SN 2022xav | 2022-11-14 | 39 | SEDM |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2022xav | 2022-11-18 | 43 | DBSP | |
SN 2022yma | 2022-11-26 | 36 | DBSP | |
SN 2022zkc | 2022-11-17 | 14 | SEDM | |
SN 2022zmb | 2022-11-18 | 15 | DBSP | |
SN 2022zmb | 2023-03-20 | 136 | LRIS | |
SN 2022zmb | 2023-04-26 | 172 | LRIS | |
SN 2022zmb | 2023-05-24 | 200 | DEIMOS | |
SN 2022zmb | 2023-06-11 | 217 | LRIS | |
SN 2022zmb | 2023-12-07 | 394 | LRIS | |
SN 2022aaad | 2022-11-27 | 18 | SEDM | |
SN 2022aaad | 2022-11-29 | 20 | ALFOSC | |
SN 2022aaad | 2022-12-04 | 25 | SEDM | |
SN 2022aaad | 2022-12-15 | 36 | SEDM | |
SN 2022aaad | 2022-12-21 | 42 | SEDM | |
SN 2022aaad | 2022-12-22 | 43 | SEDM | |
SN 2022aaad | 2022-12-29 | 50 | SEDM | |
SN 2022aaad | 2023-01-07 | 58 | SEDM | |
SN 2022aaad | 2023-01-15 | 66 | ALFOSC | |
SN 2022aaad | 2023-01-23 | 74 | SEDM | |
SN 2022aaad | 2023-02-03 | 85 | ALFOSC | |
SN 2022aaad | 2023-02-16 | 98 | SEDM | |
SN 2022aaad | 2023-02-16 | 98 | SEDM | |
SN 2022aaad | 2023-03-09 | 119 | SEDM | |
SN 2022aaad | 2023-03-24 | 134 | LRIS | |
SN 2022aaad | 2023-03-27 | 137 | SEDM | |
SN 2022aaad | 2023-04-02 | 143 | SEDM | |
SN 2022aaad | 2023-04-09 | 150 | SEDM | |
SN 2022aaad | 2023-04-12 | 153 | SEDM | |
SN 2022aaad | 2023-04-16 | 157 | ALFOSC | |
SN 2022aaad | 2023-04-20 | 161 | SEDM | |
SN 2022aaad | 2023-04-27 | 168 | DBSP | |
SN 2022aaad | 2023-04-29 | 170 | NIRES | |
SN 2022aaad | 2023-06-11 | 213 | LRIS | |
SN 2022aaad | 2023-08-18 | 281 | DBSP | |
SN 2022aaad | 2023-09-04 | 298 | SEDM | |
SN 2022aaad | 2023-09-07 | 301 | DBSP | |
SN 2022aaad | 2023-09-16 | 310 | DBSP | |
SN 2022aaad | 2023-09-21 | 315 | SEDM | |
SN 2022aaad | 2023-10-07 | 331 | SEDM | |
SN 2022aaad | 2023-10-07 | 331 | LRIS | |
SN 2022aaad | 2023-11-10 | 365 | DBSP | |
SN 2022aaad | 2023-12-07 | 392 | LRIS | |
SN 2022aaad | 2024-02-15 | 461 | LRIS | |
SN 2022aaad | 2024-03-31 | 506 | LRIS | |
SN 2022aang | 2023-03-24 | 130 | LRIS | |
SN 2022aang | 2023-12-07 | 384 | LRIS | |
SN 2022acbu | 2022-12-16 | 22 | DBSP | |
SN 2022acko | 2023-02-02 | 59 | SEDM | |
SN 2022acko | 2023-08-18 | 255 | DBSP | |
SN 2022acko | 2023-09-16 | 284 | DBSP | |
SN 2022acko | 2023-09-16 | 284 | DBSP |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2022acko | 2023-11-08 | 336 | LRIS | |
SN 2022acko | 2023-12-07 | 365 | LRIS | |
SN 2022acko | 2024-02-15 | 435 | LRIS | |
SN 2022acrl | 2022-12-19 | 18 | SEDM | |
SN 2022acri | 2022-12-18 | 16 | SEDM | |
SN 2022adtt | 2023-01-17 | 30 | LRIS | |
SN 2022adth | 2023-01-17 | 27 | LRIS | |
SN 2023cf | 2023-01-21 | 19 | SEDM | |
SN 2023axu | 2023-10-09 | 255 | SEDM | |
SN 2023blw | 2023-04-16 | 73 | DBSP | |
SN 2023blw | 2023-10-07 | 243 | LRIS | |
SN 2023bmd | 2023-03-07 | 27 | SPRAT | |
SN 2023buy | 2023-04-07 | 51 | SEDM | |
SN 2023bql | 2023-04-16 | 64 | DBSP | |
SN 1995al | 2023-03-08 | 21 | SEDM | |
SN 1995al | 2023-03-09 | 22 | SEDM | |
SN 1995al | 2023-12-07 | 294 | LRIS | |
SN 2023ghl | 2023-05-08 | 19 | ALFOSC | |
SN 2023gss | 2023-06-10 | 45 | SEDM | |
SN 2023gxq | 2023-05-24 | 26 | DEIMOS | |
SN 2023gxq | 2023-05-28 | 30 | DBSP | |
SN 2023gxq | 2023-06-11 | 44 | LRIS | |
SN 2023gxq | 2023-06-12 | 45 | LRIS | |
SN 2023hcp | 2023-05-15 | 16 | SEDM | |
SN 2023hlf | 2023-05-01 | 12 | SEDM | |
SN 2023ijd | 2023-05-28 | 14 | SEDM | |
SN 2023ijd | 2023-12-03 | 201 | SEDM | |
SN 2023ijd | 2023-12-03 | 201 | SEDM | |
SN 2023ijd | 2023-12-06 | 204 | SEDM | |
SN 2023hzt | 2023-05-28 | 20 | DBSP | |
SN 2023hzt | 2023-06-14 | 36 | SEDM | |
SN 2023jid | 2023-05-28 | 95 | DBSP | |
SN 2023kne | 2023-06-20 | 19 | DBSP | |
SN 2023kzz | 2023-08-18 | 65 | DBSP | |
SN 2023rpu | 2023-09-28 | 108 | SEDM | |
SN 2023lzn | 2023-07-10 | 10 | SEDM | |
SN 2023mpj | 2023-07-11 | 15 | DBSP | |
SN 2023nca | 2023-07-18 | 13 | LRIS | |
SN 2023nca | 2023-07-23 | 18 | LRIS | |
SN 2023ngy | 2023-07-28 | 13 | SEDM | |
SN 2023ngy | 2023-08-01 | 17 | DBSP | |
SN 2023ngy | 2023-08-03 | 19 | SEDM | |
SN 2023nlu | 2023-08-04 | 17 | SEDM | |
SN 2023vhb | 2023-11-14 | 80 | SEDM | |
SN 2023rix | 2023-09-17 | 12 | SEDM | |
SN 2023rtq | 2023-09-24 | 17 | SEDM | |
SN 2023rtq | 2023-10-07 | 30 | LRIS | |
SN 2023rvo | 2023-09-19 | 13 | SEDM | |
SN 2023rvo | 2023-10-05 | 28 | DBSP | |
SN 2023ucx | 2023-10-07 | 13 | LRIS | |
SN 2023udb | 2023-10-15 | 13 | Deveny+LMI |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2023udb | 2023-10-17 | 15 | LRIS | |
SN 2023twg | 2023-10-17 | 19 | SEDM | |
SN 2023vcj | 2023-12-09 | 57 | SEDM | |
SN 2023vcj | 2023-12-10 | 58 | ALFOSC | |
SN 2023vcj | 2023-12-28 | 76 | SEDM | |
SN 2023vcj | 2023-12-29 | 77 | SEDM | |
SN 2023vcj | 2023-12-30 | 78 | ALFOSC | |
SN 2023vcj | 2024-01-06 | 84 | SEDM | |
SN 2023vcj | 2024-01-06 | 84 | SEDM | |
SN 2023vcj | 2024-01-08 | 86 | KAST | |
SN 2023vcj | 2024-01-19 | 97 | ALFOSC | |
SN 2023vcj | 2024-01-26 | 104 | SEDM | |
SN 2023vcj | 2024-02-11 | 120 | SEDM | |
SN 2023vcj | 2024-03-09 | 146 | ALFOSC | |
SN 2023vcj | 2024-03-14 | 151 | ALFOSC | |
SN 2023vcj | 2024-03-30 | 166 | ALFOSC | |
SN 2023vog | 2023-11-01 | 11 | SEDM | |
SN 2023vog | 2023-11-13 | 23 | SEDM | |
SN 2023way | 2023-11-03 | 10 | SEDM | |
SN 2023way | 2023-11-05 | 12 | SEDM | |
SN 2023wcr | 2023-12-03 | 40 | DBSP | |
SN 2023wcr | 2024-01-13 | 81 | DBSP | |
SN 2023wcr | 2024-01-13 | 81 | DBSP | |
SN 2023wcr | 2024-02-15 | 114 | LRIS | |
SN 2023wcr | 2024-03-15 | 143 | ALFOSC | |
SN 2023wcr | 2024-03-19 | 147 | SEDM | |
SN 2023wcr | 2024-03-20 | 148 | DBSP | |
SN 2023wcr | 2024-04-03 | 162 | SEDM | |
SN 2023wcr | 2024-04-12 | 171 | DBSP | |
SN 2023wcr | 2024-05-01 | 190 | DBSP | |
SN 2023wcr | 2024-07-11 | 260 | DBSP | |
SN 2023xvo | 2023-11-21 | 34 | SEDM | |
SN 2023xvo | 2023-12-03 | 45 | DBSP | |
SN 2023zcu | 2024-01-10 | 31 | SEDM | |
SN 2023zcu | 2024-11-11 | 335 | DBSP | |
SN 2023acbr | 2024-02-15 | 53 | LRIS | |
SN 2024V | 2024-02-15 | 53 | LRIS | |
SN 2024ov | 2024-01-13 | 30 | DBSP | |
SN 2024ov | 2024-03-31 | 107 | LRIS | |
SN 2024ov | 2024-05-08 | 144 | LRIS | |
SN 2024ov | 2024-11-30 | 345 | LRIS | |
SN 2024wp | 2024-04-12 | 90 | DBSP | |
SN 2024wp | 2024-05-02 | 110 | DBSP |
Source | Date | Phase | Inst. | H I 6563 |
(days) | (km s-1) | |||
SN 2024wp | 2024-07-09 | 177 | LRIS | |
SN 2024aul | 2024-02-27 | 36 | SPRAT | |
AT 2024bzq | 2024-02-23 | 16 | SEDM | |
SN 2024btx | 2024-05-08 | 90 | LRIS | |
SN 2024atk | 2024-04-10 | 66 | SEDM | |
SN 2024chx | 2024-04-03 | 50 | SPRAT | |
SN 2024daa | 2024-03-20 | 33 | DBSP | |
SN 2024dhi | 2024-04-12 | 50 | DBSP | |
SN 2024egd | 2024-03-18 | 15 | ALFOSC | |
SN 2024ees | 2024-04-02 | 25 | SPRAT | |
SN 2024etq | 2024-03-31 | 19 | LRIS | |
SN 2024epy | 2024-03-31 | 15 | LRIS | |
SN 2024faf | 2024-04-12 | 26 | DBSP | |
SN 2024grw | 2024-04-27 | 11 | SPRAT | |
SN 2024grw | 2024-05-06 | 20 | SEDM | |
SN 2024hme | 2024-05-07 | 18 | DBSP | |
SN 2024izq | 2024-05-31 | 20 | DBSP | |
SN 2024jlf | 2024-06-07 | 11 | Binospec | |
SN 2024jlf | 2024-06-07 | 11 | SEDM | |
SN 2024jlf | 2024-06-10 | 14 | SEDM | |
SN 2024jlf | 2024-06-10 | 14 | SEDM | |
SN 2024jlf | 2024-06-12 | 16 | SEDM | |
SN 2024jlf | 2024-06-18 | 22 | SEDM | |
SN 2024jlf | 2024-06-19 | 23 | SEDM | |
SN 2024jlf | 2024-06-29 | 33 | LRIS | |
SN 2024jlf | 2024-07-04 | 37 | SEDM | |
SN 2024jlf | 2024-07-06 | 39 | SEDM | |
SN 2024jlf | 2024-07-16 | 49 | SEDM | |
SN 2024jlf | 2024-07-27 | 60 | SEDM | |
SN 2024jlf | 2024-08-01 | 65 | ALFOSC | |
SN 2024jlf | 2024-08-14 | 78 | SEDM | |
SN 2024jlf | 2024-08-26 | 90 | SEDM | |
SN 2024jlf | 2024-08-30 | 94 | KCWI | |
SN 2024jxm | 2024-06-15 | 16 | DBSP | |
SN 2024jxm | 2024-08-28 | 89 | DBSP | |
SN 2024jxm | 2024-09-04 | 96 | SEDM | |
SN 2024jxm | 2024-09-05 | 97 | SEDM | |
SN 2024jxm | 2024-10-09 | 130 | LRIS | |
SN 2024jxm | 2024-11-30 | 181 | LRIS | |
SN 2024jxm | 2025-01-26 | 237 | LRIS | |
SN 2024ldu | 2024-06-18 | 12 | SEDM | |
SN 2024ldu | 2024-07-09 | 33 | LRIS | |
SN 2024mxq | 2024-07-09 | 19 | DBSP |